### v1

# First-passage properties of the jump process with a drift. The general case

Abstract.
We study the first-passage properties of a jump process with constant drift where jump amplitudes and inter-arrival times follow arbitrary light-tailed distributions with smooth densities. Using a mapping to an effective discrete-time random walk, we identify three regimes determined by the drift strength: survival (weak drift), absorption (strong drift), and critical. We derive explicit expressions for exponential decay rates in the survival and absorption regimes, and characterize algebraic decay at the critical point. We also obtain asymptotic behavior of the mean first-passage time, number of jumps, and their variances for processes starting either close to the origin or far from it.
1. Introduction
Imagine a bakery with a single counter where a cashier serves customers who arrive at random times, each requiring a varying amount of time to complete their order. The cashier works at a consistent pace, attending to one customer at a time in the order they arrive, with no overtaking allowed. This setup, that we have so frivolously described, actually represents a fundamental model in queuing theory. Specifically, if both the inter-arrival times and the amount of time it takes to serve a customer are independent identically distributed random variables, this process is usually referred to as a G/G/1 queue (see [1] for a pedagogical introduction).
A simple observation is that if many customers arrive with complex orders, the queue will grow indefinitely. In contrast, if few customers arrive or orders are simple to fulfill, the cashier will remain idle most of the time. So it is natural to wonder whether the queue depletes and, if so, how much time does it take? This is essentially the question we address in the present paper.
Figure 1. An example of the trajectory. Starting at $X_{0}$ the process instantaneously undergoes a jump $M_{1}$, then moves toward the origin with constant velocity $\alpha$ for the time $t_{1}$ when the next jump $M_{2}$ occurs. This pattern continues until the process crosses the origin at time $\tau$ after $n$ jumps (here $n=6$). The inter-jump intervals $\{t_{1},\ldots,t_{6}\}$ are i.i.d. random variables drawn from $p(t)$, and the jump amplitudes $\{M_{1},\ldots,M_{6}\}$ are also i.i.d. random variables following $q(M)$.
To formulate the problem in a more formal manner, we consider a jump process with a drift, i.e., a stochastic process $X(t)$ evolving on the real line through a combination of deterministic negative linear drift and random positive jumps (see Fig. 1). In the <<bakery>> picture, $X(t)$ is the total amount of requests piled in the queue, the linear decay segments correspond to the service of customers, while the jumps represent the sporadic arrivals of newcomers. The dynamics of $X(t)$ unfold as follows:
  1. (1)
    The process starts at an initial position $X_{0}$, and immediately experiences a jump of random amplitude $M_{1}$. It then drifts toward the origin at constant speed $\alpha$ for a random time interval $t_{1}$, after which the next jump $M_{2}$ occurs.
  2. (2)
    This cycle repeats until the first-passage event, i.e., until the process crosses the origin to the negative side for the first time (the queue is depleted).
  3. (3)
    The time intervals between successive jumps, $t_{i}$, are independent and identically distributed (i.i.d.), drawn from a distribution with density $p(t)$.
  4. (4)
    The jump amplitudes $M_{i}$ are also i.i.d., sampled from a distribution $q(M)$.
The model is controlled by four parameters: the initial position $X_{0}\geq 0$, the drift strength $\alpha>0$, and the two probability distributions $p(t)$ and $q(M)$. Formally, the process can be written as
$X(t)=X_{0}-\alpha t+\sum_{j=1}^{n(t)}M_{j},$ (1)
where $n(t)$ is a renewal process describing the number of jumps that have occurred up to time $t$.
Our aim is to study the first-passage properties of the process (1). Specifically, we focus on two observables: the first-passage time $\tau$ and the number of jumps $n$ that occur before the first-passage event. Clearly, these two quantities are random variables, which are in general correlated.
[…]
In [36] we proposed an alternative approach that does not rely on the renewal equation. Specifically, we constructed a mapping of the original continuous-time process onto an effective discrete-time random walk. This mapping allowed us to obtain a closed-form representation of the Laplace transform of the joint probability distribution of $\tau$ and $n$, by applying the Pollaczek-Spitzer formula [37, 38, 39], a powerful result in discrete random walk theory. More precisely, we used a generalization [40] to the asymmetric random walk of the formula obtained by Ivanov in [41]. This approach was then successfully applied to two exactly solvable cases with Poisson arrivals, where explicit solutions could be obtained and compared against standard renewal equation techniques.
The present work tackles a significantly more challenging general case, where explicit solutions are no longer available. Here, we consider a rather general class of distributions requiring that $p(t)$ and $q(M)$ be light-tailed with smooth probability densities, thereby moving far beyond the assumption of Poisson arrivals. In this general case, the closed-form integral representations of the relevant Laplace transforms are still valid, but extracting the asymptotic behavior now requires analytic continuation techniques and singularity analysis. In this work we provide such analysis and derive several exact asymptotic results for this general setting.
The main results concern the behavior of the survival probabilities, i.e., the probability that the process stays positive up to time $\tau$ (or up to the $n$-th jump). We show that there are three distinct regimes determined by the drift strength. In the survival regime (weak drift), the process has a finite probability of never crossing the origin, and survival probabilities decay exponentially to a positive constant. In the absorption regime (strong drift), the first passage event occurs with certainty, and survival probabilities decay exponentially to zero. We compute these exponential decay rates explicitly for arbitrary light-tailed distributions with smooth densities. At the critical point separating these regimes, exponential decay is replaced by algebraic decay. Such behavior mirrors that obtained in [36] for two exactly solvable cases.
Beyond the decay rates of the survival probabilities, we obtain exact asymptotic expansions for several quantities as $X_{0}\to 0$ and $X_{0}\to\infty$. For the absorption regime, we derive expansions for the mean and variance of both $\tau$ and $n$. For the survival regime, we characterize the asymptotic behavior of the infinite-time survival probability. At the critical point, we establish the algebraic decay and derive asymptotic expansions for the amplitudes of the power-law behavior.
The paper is organized as follows: In Section 2 we define the model more formally and recall the representation for the triple Laplace transform (with respect to $\tau$, $n$ and $X_{0}$) of the joint probability distribution of $\tau$ and $n$ obtained in [36]. In Section 3 we study the properties of this transform in the $\lambda$-plane ($\lambda$ being the Laplace dual of $X_{0}$) and perform analytic continuation. Sections 4, 5 and 6 analyze the asymptotic behavior of the survival probabilities: at infinite time, up to the $n$-th jump and up to time $\tau$, respectively. Section 7 concerns the asymptotic expansions of the means and variances of $\tau$ and $n$ in the absorption regime. Finally, we conclude in Section 8.
2. The model and the main results
In this paper we consider the process (1) in a rather general setup. That is, we do not specify the distributions $p(t)$ and $q(M)$ and instead only require them to have two properties:
  1. (1)
    All moments of both $M$ and $t$ are finite (both distributions are light-tailed).
  2. (2)
    The densities $p(t)$ and $q(M)$ are continuously differentiable functions, supported either on a finite interval $0,\ell]$ or on the positive half-line $0,\infty)$.
We start by fixing notation and defining the key quantities. We denote by $\mathbb{P}[\tau,n\,|\,X_{0}]$ the joint probability distribution of the first-passage time $\tau$ and the number of jumps $n$ for the process started at $X_{0}$. Our analysis focuses on three survival probabilities:
  • The probability that the process remains positive until time $\tau$:
    $S_{T}(\tau\,|\,X_{0})\equiv 1-\int_{0}^{\tau}\mathrm{d}\bar{\tau}\sum_{k=0}^{\infty}\mathbb{P}[\bar{\tau},k\,|\,X_{0}].$ (2)
  • The probability that the process remains positive for at least $n$ jumps:
    $S_{N}(n\,|\,X_{0})\equiv 1-\int_{0}^{\infty}\mathrm{d}\bar{\tau}\sum_{k=0}^{n}\mathbb{P}[\bar{\tau},k\,|\,X_{0}].$ (3)
  • The probability that the process remains positive indefinitely:
    $S_{\infty}(X_{0})\equiv 1-\int_{0}^{\infty}\mathrm{d}\bar{\tau}\sum_{k=0}^{\infty}\mathbb{P}[\bar{\tau},k\,|\,X_{0}].$ (4)
Before presenting exact results, let us briefly provide some intuition about the expected behavior. If the drift is strong, it pulls the process toward the origin, so the first-passage event certainly happens and $S_{\infty}(X_{0})=0$. If the drift is weak, the jumps dominate and the process can escape to infinity, so there is a finite probability that first-passage never occurs and $S_{\infty}(X_{0})>0$.
The transition between these regimes can be understood by examining the average displacement per cycle. Each cycle consists of a jump of size $M$ followed by drift of $\alpha t$ toward the origin. The sign of the average displacement determines the regime:
$\langle M-\alpha t\rangle>0$ $:\qquad S_{\infty}(X_{0})>0,$ (5)
$\langle M-\alpha t\rangle<0$ $:\qquad S_{\infty}(X_{0})=0,$
where $\langle\cdots\rangle$ denotes averaging with respect to $p(t)$ and $q(M)$. This simple observation gives the critical drift velocity:
$\alpha_{c}=\frac{\langle M\rangle}{\langle t\rangle}.$ (6)
We treat the survival probability at infinite time as an <<order parameter>> separating the survival regime, where $S_{\infty}(X_{0})>0$, from the absorption regime, where $S_{\infty}(X_{0})=0$, with $\alpha$ being the <<control parameter>>. The case in which $\alpha=\alpha_{c}$ we refer to as the critical point (see Fig. 2 for a schematic representation).
Figure 2. Schematic representation of the survival probability $S_{\infty}(X_{0})$ as a function of the drift strength $\alpha$ for fixed initial position $X_{0}$. Below the critical value $\alpha_{c}$, the system is in the survival regime with $S_{\infty}(X_{0})>0$. The critical point occurs at $\alpha=\alpha_{c}$. Above $\alpha_{c}$, the system is in the absorption regime with $S_{\infty}(X_{0})=0$.
The key ingredient in the analysis is the closed-form representation of the Laplace transform $\hat{Q}(\rho,s\,|\,\lambda)$ of the joint probability distribution $\mathbb{P}[\tau,n\,|\,X_{0}]$,
$\hat{Q}(\rho,s\,|\,\lambda)\equiv\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\int_{0}^{\infty}{\mathrm{d}\tau}\,e^{-\rho\tau}\,\sum_{n=0}^{\infty}s^{n}\,\mathbb{P}[\tau,n\,|\,X_{0}].$ (7)
[…]
$\hat{Q}(\rho,s\,|\,\lambda)=\frac{1}{\lambda+\frac{\rho}{\alpha}}-\frac{1-s\,c(\rho)}{\lambda+\frac{\rho}{\alpha}}\;\phi^{-}\!\left(\lambda+\frac{\rho}{\alpha};\rho,s\right)\phi^{+}\!\left(0;\rho,s\right),$ (8)
where
$\phi^{\pm}(\lambda;\rho,s)\equiv\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\;\frac{1}{\lambda\pm\mathrm{i}k}\log\left[1-s\,c(\rho)F(k;\rho)\right]\right],$ (9)
and
$c(\rho)\equiv\int_{0}^{\infty}\mathrm{d}M\,e^{-\rho\frac{M}{\alpha}}q(M).$ (10)
The function $F(k;\rho)$ entering (9) is a Fourier transform of the probability distribution of the steps in the effective random walk. In terms of the original probability distributions $p(t)$ and $q(M)$ it reads
$F(k;\rho)\equiv\frac{1}{c(\rho)}\int_{0}^{\infty}\mathrm{d}M\,e^{-\rho\frac{M}{\alpha}+\mathrm{i}kM}q(M)\int_{0}^{\infty}\mathrm{d}t\,e^{-\mathrm{i}\alpha kt}p(t),$ (11)
or alternatively
$F(k;\rho)=\frac{1}{c(\rho)}\left\langle e^{-\rho\frac{M}{\alpha}+\mathrm{i}k(M-\alpha t)}\right\rangle.$ (12)
The joint probability distribution $\mathbb{P}[\tau,n\,|\,X_{0}]$ could, in principle, be found by inverting the transform in (7). This is, however, a very thorny path and there is little reason to hope that it can be traversed for arbitrary distributions. Instead we study the analytic structure of the triple Laplace transform $\hat{Q}(\rho,s\,|\,\lambda)$ and obtain the asymptotic results. The main caveat is that the parameters $\lambda$, $\rho$ and $s$ originate from the Laplace transforms; hence the representation (9) is valid for $\lambda>0$, $\rho>0$, and $s\in(0,1)$. However, extracting the asymptotic behavior of the probabilities requires extending this representation to larger domains through proper analytic continuation.
[…]
Finally, we make two remarks. First, we note that alternative representations for survival probabilities of discrete-time random walks have recently been obtained using Riemann-Hilbert techniques [57], providing a complementary approach to the Pollaczek-Spitzer formula. Second, one of the two exactly solvable cases in [36] takes $q(M)$ to be a delta function; hence, the smoothness requirement is not fulfilled, but the detailed computation using (8) can still be performed. The reason is that this requirement is mostly technical and can be lifted entirely for a significant portion of the derivations. In addition, the results can readily be extended to cases where only the first three moments of $M$ and $t$ are finite. Keeping this level of precision, however, comes at the cost of significantly more technical detail without offering much additional insight. For this reason, we choose to sacrifice a small degree of generality in favor of clarity, and throughout the paper we always assume that both distributions $p(t)$ and $q(M)$ have smooth densities and that all their moments are finite.
2.1. The main results
By leveraging the representation (8) of the triple Laplace transform, we derive asymptotic results valid for arbitrary probability distributions $p(t)$ and $q(M)$. The key advantage is that all expressions are written in terms of the Fourier transform $F(k;\rho)$ of the effective random walk (11).
Survival regime
When the drift is weak ($\alpha<\alpha_{c}$), there is a finite probability that the process remains positive indefinitely, $S_{\infty}(X_{0})>0$. The key result here is explicit expressions for the exponential decay rates that govern the approach to this asymptotic value.
Both survival probabilities $S_{N}(n\,|\,X_{0})$ and $S_{T}(\tau\,|\,X_{0})$ approach their infinite-time limit $S_{\infty}(X_{0})$ exponentially fast:
$\alpha<\alpha_{c}:\quad\left\{\begin{aligned} S_{N}(n\,|\,X_{0})-S_{\infty}(X_{0})&\underset{n\to\infty}{\asymp}\exp\left[-\frac{n}{\xi_{n}(\alpha)}\right],\\ S_{T}(\tau\,|\,X_{0})-S_{\infty}(X_{0})&\underset{\tau\to\infty}{\asymp}\exp\left[-\frac{\tau}{\xi_{\tau}(\alpha)}\right].\end{aligned}\right$. (13)
Crucially, the decay rates are given by explicit, computable expressions:
$\frac{1}{\xi_{n}(\alpha)}=-\log\left[\min_{k_{I}\in(\zeta_{-},\zeta_{+})}F(\mathrm{i}k_{I};0)\right],\qquad\frac{1}{\xi_{\tau}(\alpha)}=-\rho_{*},$ (14)
where $\rho_{*}<0$ is determined by the condition
$\frac{1}{c(\rho_{*})}=\min_{k_{I}\in(\zeta_{-},\zeta_{+})}F(\mathrm{i}k_{I};\rho_{*}).$ (15)
The minimization is performed over purely imaginary arguments within the region where $\mathrm{Im}(k)\in(\zeta_{-},\zeta_{+})$}, which is the maximal strip in the complex $k$-plane where the Fourier transform $F(k;\rho)$ defined in (11) is analytic (see Section 3.1 for details).
Additionally, we obtain the asymptotic behavior of $S_{\infty}(X_{0})$ itself. If the process starts close to the origin, then
$S_{\infty}(X_{0})\underset{X_{0}\to 0}{\sim}\sqrt{\frac{2\langle M-\alpha t\rangle^{2}}{\langle(M-\alpha t)^{2}\rangle}}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{X_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\log\left|1-F(k;0)\right|\right).$ (16)
For initial positions far from the origin, we obtain exponential behavior
$1-S_{\infty}(X_{0})\underset{X_{0}\to\infty}{\asymp}e^{-RX_{0}},$ (17)
where $R>0$ is the smallest positive solution of $F(\mathrm{i}R;0)=1$, analogous to the Lundberg exponent in ruin theory.
Critical point
As the drift approaches the critical value, the rates (14) diverge:
$\xi_{n}(\alpha)\underset{\alpha\to\alpha_{c}}{\sim}\frac{2}{\langle t\rangle^{2}}\frac{\langle(M-\alpha_{c}t)^{2}\rangle}{(\alpha-\alpha_{c})^{2}},\qquad\xi_{\tau}(\alpha)\underset{\alpha\to\alpha_{c}}{\sim}\frac{2}{\langle t\rangle}\frac{\langle(M-\alpha_{c}t)^{2}\rangle}{(\alpha-\alpha_{c})^{2}}.$ (18)
This leads to an algebraic decay of the survival probabilities at the critical point $\alpha=\alpha_{c}$:
$\alpha=\alpha_{c}:\quad\left\{\begin{aligned} &S_{N}(n\,|\,X_{0})\underset{n\to\infty}{\sim}\sqrt{\frac{1}{\pi n}}U(X_{0}),\\ &S_{T}(\tau\,|\,X_{0})\underset{\tau\to\infty}{\sim}\sqrt{\frac{\langle t\rangle}{\pi\tau}}U(X_{0}),\end{aligned}\right$. (19)
where the function $U(X_{0})$ is defined through its Laplace transform in (109). This function admits the following asymptotic expansions:
$U(X_{0})\underset{X_{0}\to\infty}{\sim}U_{1}X_{0}+U_{0},\qquad U(X_{0})\underset{X_{0}\to 0}{\sim}u_{0}+u_{1}X_{0},$ (20)
with explicit coefficients $U_{0,1}$ and $u_{0,1}$ given in (110) and (111).
In addition, we show that at criticality, there exists a Brownian scaling limit. When $n\to\infty$ and $X_{0}\to\infty$ with $X_{0}/\sqrt{n}$ fixed, we find the scaling behavior
$S_{N}(n\,|\,X_{0})\underset{n\to\infty}{\sim}\text{erf}\left(\frac{X_{0}}{\sqrt{2n\langle(M-\alpha_{c}t)^{2}\rangle}}\right),\quad\frac{X_{0}}{\sqrt{n}}\text{\penalty 10000\ --- fixed}.$ (21)
Similarly, when $\tau\to\infty$ and $X_{0}\to\infty$ with $X_{0}/\sqrt{\tau}$ fixed:
$S_{T}(\tau\,|\,X_{0})\underset{\tau\to\infty}{\sim}\text{erf}\left(\frac{X_{0}}{\sqrt{2\frac{\tau}{\langle t\rangle}\langle(M-\alpha_{c}t)^{2}\rangle}}\right),\quad\frac{X_{0}}{\sqrt{\tau}}\text{\penalty 10000\ --- fixed}.$ (22)
Absorption regime
When the drift is strong ($\alpha>\alpha_{c}$), the process eventually crosses the origin with certainty. The survival probabilities again decay exponentially:
$\alpha>\alpha_{c}:\quad\left\{\begin{aligned} S_{N}(n\,|\,X_{0})&\underset{n\to\infty}{\asymp}\exp\left[-\frac{n}{\xi_{n}(\alpha)}\right],\\ S_{T}(\tau\,|\,X_{0})&\underset{\tau\to\infty}{\asymp}\exp\left[-\frac{\tau}{\xi_{\tau}(\alpha)}\right].\end{aligned}\right$. (23)
The rates given by the same expressions as in the survival regime (14).
In this regime the moments of $\tau$ and $n$ are finite, and their Laplace transform with respect to $X_{0}$ can be found from (8). Here we obtain the first two terms in the asymptotic expansions as $X_{0}\to\infty$ and $X_{0}\to 0$ for the mean and variance of both $n$ and $\tau$:
$\mathbb{E}[n\,|\,X_{0}]\underset{X_{0}\to\infty}{\sim}X_{0}A_{1}+A_{0},\qquad\mathbb{E}[n\,|\,X_{0}]\underset{X_{0}\to 0}{\sim}a_{0}+a_{1}X_{0},$ (24)
$\mathbb{E}[\tau\,|\,X_{0}]\underset{X_{0}\to\infty}{\sim}X_{0}\tilde{A}_{1}+\tilde{A}_{0},\qquad\mathbb{E}[\tau\,|\,X_{0}]\underset{X_{0}\to 0}{\sim}\tilde{a}_{0}+\tilde{a}_{1}X_{0},$ (25)
$\mathrm{Var}[n\,|\,X_{0}]\underset{X_{0}\to\infty}{\sim}X_{0}B_{1}+B_{0},\qquad\mathrm{Var}[n\,|\,X_{0}]\underset{X_{0}\to 0}{\sim}b_{0}+b_{1}X_{0},$ (26)
$\mathrm{Var}[\tau\,|\,X_{0}]\underset{X_{0}\to\infty}{\sim}X_{0}C_{1}+C_{0},\qquad\mathrm{Var}[\tau\,|\,X_{0}]\underset{X_{0}\to 0}{\sim}c_{0}+c_{1}X_{0}.$ (27)
All coefficients are expressed as integrals involving $F(k;\rho)$, with complete expressions provided in Section 7.
Let us conclude the presentation of the results by providing some motivation for constructing expansions beyond the leading order. While subleading coefficients might appear as technical details, in some cases they carry significant physical meaning. A notable example is the constant $U_{0}$ in the critical regime expansion (110), which in the symmetric random walk case is closely related to the Milne extrapolation length in radiative transport theory [58] and trapping problem for Rayleigh flights [59] (see [51] for a more comprehensive overview). Similar constants also find applications in packing problems in computer science [60, 61]. Since these corrections are nontrivial to obtain, we believe it is worthwhile to present here a systematic derivation based on Mellin transform techniques, which provides a robust framework for extracting such asymptotic expansions.
3. Analytic structure of $\phi^{\pm}(\lambda;\rho,s)$ in the $\lambda$-plane
The parameters $\lambda$, $s$, and $\rho$ in (9) arise from Laplace transforms. Hence, the representation is initially valid for $\lambda>0$, $s\in(0,1)$, and $\rho>0$. The asymptotic analysis of first-passage quantities, however, requires an analytic continuation of $\phi^{\pm}(\lambda;\rho,s)$ to broader domains. As we will shortly show, this continuation is not as simple as extending (9) to, say, $\lambda<0$; it demands much more careful consideration.
In this section, we treat $\phi^{\pm}(\lambda;\rho,s)$ as functions of $\lambda$ and extend (9) to the complex $\lambda$-plane. We begin by establishing several key properties of the Fourier transform $F(k;\rho)$ given by (11), which serves as a crucial elementary building block. Next, we consider a simple yet illustrative example to show why the analytic continuation of the integrals such as (9) differs from naively applying the definition for $\lambda<0$. Guided by this example, we construct the analytic continuations $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ of $\phi^{\pm}(\lambda;\rho,s)$ to the $\lambda$-plane. In this section we assume that $s<1$ and $\rho>0$; continuations in the $s$- and $\rho$-planes are addressed in the next sections.
3.1. Properties of $F(k;\rho)$
The key ingredient in $\phi^{\pm}(\lambda;\rho,s)$ is the Fourier transform $F(k;\rho)$ of the effective random walk. Our analysis therefore relies on several properties of $F(k;\rho)$. In particular, we use its asymptotic expansions for real $k$ in the limits $k\to 0$ and $k\to\infty$, as well as its analytic structure in the complex $k$-plane.
Expansions for real $k$
The expansion of $F(k;\rho)$ as $k\to 0$ is straightforward. Since all moments of $p(t)$ and $q(M)$ are finite, so are the moments of the effective random walk, and thus the expansion reads:
$F(k;\rho)\underset{k\to 0}{=}1+\mathrm{i}\mu_{1}(\rho)k-\mu_{2}(\rho)\frac{k^{2}}{2}-\mu_{3}(\rho)\frac{\mathrm{i}k^{3}}{6}+O\left(k^{4}\right),$ (28)
where $\mu_{\ell}(\rho)$ denotes the $\ell$-th moment of the effective random walk,
$\mu_{\ell}(\rho)\equiv\frac{1}{c(\rho)}\int_{0}^{\infty}\mathrm{d}t\,\int_{0}^{\infty}\mathrm{d}M\,e^{-\rho\frac{M}{\alpha}}\left(M-\alpha t\right)^{\ell}q(M)\,p(t),$ (29)
or, equivalently,
$\mu_{\ell}(\rho)=\frac{1}{c(\rho)}\left\langle e^{-\rho\frac{M}{\alpha}}\left(M-\alpha t\right)^{\ell}\right\rangle.$ (30)
For the expansion as $k\to\infty$, note that $F(k;\rho)$ is a product of two highly oscillatory integrals, each decaying at large $k$. Consider the second integral, assuming that $p(t)$ is supported on $0,\infty)$. Integration by parts yields
$\int_{0}^{\infty}\mathrm{d}t\,p(t)e^{-\mathrm{i}k\alpha t}=\left.-p(t)\frac{e^{-\mathrm{i}k\alpha t}}{\mathrm{i}k\alpha}\right|_{0}^{\infty}+\frac{1}{\mathrm{i}k\alpha}\int_{0}^{\infty}\mathrm{d}t\,p^{\prime}(t)e^{-\mathrm{i}\alpha kt},$ (31)
where the boundary term at infinity vanishes due to the decay of $p(t)$. The leading-order behavior is thus determined by the contribution at $t=0$. Since $p(t)$ is continuously differentiable and light-tailed, the remaining integral is subdominant. Hence
$\int_{0}^{\infty}\mathrm{d}t\,p(t)e^{-\mathrm{i}k\alpha t}\underset{k\to\infty}{=}O\left(\frac{1}{k}\right).$ (32)
If the probability density is supported on a finite interval instead of the positive real axis, the result includes oscillatory terms, but the absolute value remains of order $1/k$. Repeating the same argument for the integral over $M$ in (11), we find that it is also $O(1/k)$ as $k\to\infty$. Thus,
$F(k;\rho)\underset{k\to\infty}{=}O\left(\frac{1}{k^{2}}\right).$ (33)
Analytic properties in the $k$-plane
The next step is to analytically continue $F(k;\rho)$ into the complex $k$-plane. Since the distributions $p(t)$ and $q(M)$ are light-tailed, the function defined by (11) is analytic in the complex strip $\zeta_{-}(\rho)<\,\mathrm{Im}(k)<\zeta_{+}$ (see Fig. 3). The bounds of this strip are determined by the decay rates of $p(t)$ and $q(M)$ as $t\to\infty$ and $M\to\infty$. Specifically, $\zeta_{+}>0$ and $\zeta_{-}(\rho)<0$ are $k$-independent quantities such that
$p(t)\underset{t\to\infty}{\lesssim}\exp\left[-\zeta_{+}\,\alpha t\right],\qquad q(M)\underset{M\to\infty}{\lesssim}\exp\left[\left(\zeta_{-}(\rho)+\frac{\rho}{\alpha}\right)M\right].$ (34)
The bounds in (34) may be equivalently defined as the edges of the region where the integral (11) converges absolutely.
In what follows, we write $\zeta_{-}\equiv\zeta_{-}(\rho)$ omitting the explicit $\rho$-dependence to simplify notation. In cases where the probability distributions decay faster than exponential (e.g., if $p(t)$ or $q(M)$ is half-Gaussian), we set $\zeta_{\pm}=\pm\infty$. Furthermore, when referring to $F(k;\rho)$, we will always assume that $k$ lies within the strip of convergence, i.e., $k=k_{R}+\mathrm{i}k_{I}$ with $k_{R}\in\mathbb{R}$ and $k_{I}\in(\zeta_{-},\zeta_{+})$.
Figure 3. The structure of the Fourier transform $F(k;\rho)$ in the complex $k$-plane. The function is analytic in the horizontal strip $\mathrm{Im}(k)\in(\zeta_{-},\zeta_{+})$, where $\zeta_{\pm}$ are constants determined by the decay rates of the probability densities $p(t)$ and $q(M)$, as given in (34).
The first simple observation about $F(k;\rho)$ is that within the strip of convergence, we have
$\left|F(k_{R}+\mathrm{i}k_{I};\rho)\right|<\left|F(\mathrm{i}k_{I};\rho)\right|,\qquad k_{R}\neq 0,\quad k_{I}\in(\zeta_{-},\zeta_{+}).$ (35)
This inequality arises because the real part $k_{R}$ introduces oscillations in the integral (11), reducing its absolute value compared to the case where $k$ is purely imaginary.
Since the absolute value of $F(k;\rho)$ is bounded by its values along the imaginary axis, let us examine $F(\mathrm{i}k_{I};\rho)$ more closely. From (11), it immediately follows that
$F(\mathrm{i}k_{I};\rho)=\frac{1}{c(\rho)}\int_{0}^{\infty}\mathrm{d}M\,q(M)\int_{0}^{\infty}\mathrm{d}t\,p(t)\;e^{-\rho\frac{M}{\alpha}-k_{I}(M-\alpha t)},$ (36)
and hence
$F(\mathrm{i}k_{I};\rho)\Big|_{k_{I}=0}=1,\qquad\frac{\partial}{\partial k_{I}}F(\mathrm{i}k_{I};\rho)\Big|_{k_{I}=0}=-\mu_{1}(\rho).$ (37)
Furthermore, for the second derivative we have
$\frac{\partial^{2}}{\partial k_{I}^{2}}F(\mathrm{i}k_{I};\rho)=\frac{1}{c(\rho)}\int_{0}^{\infty}\mathrm{d}M\,q(M)\int_{0}^{\infty}\mathrm{d}t\,p(t)\,e^{-\rho\frac{M}{\alpha}-k_{I}(M-\alpha t)}(M-\alpha t)^{2}.$ (38)
All factors in the integrand are positive; hence $F(\mathrm{i}k_{I};\rho)$ is strictly convex. This implies a unique minimum, located at some point $k_{I}=\zeta_{*}(\rho)$. The position of $\zeta_{*}(\rho)$ depends on the sign of $\mu_{1}(\rho)$: it lies to the left of the origin if $\mu_{1}(\rho)>0$ and to the right if $\mu_{1}(\rho)<0$ (see Fig. 4 for a schematic representation).
Moreover, we can conclude that, if $sc(\rho)<1$, which is indeed satisfied for $s\in(0,1)$ and $\rho>0$, the equation
$1-s\,c(\rho)\,F(\mathrm{i}\zeta;\rho)=0$ (39)
admits two real solutions: a negative one $\zeta_{1}(\rho,s)$ and a positive one $\zeta_{2}(\rho,s)$. To simplify notation, we will often omit the explicit $s$- and $\rho$-dependence and write $\zeta_{1,2}\equiv\zeta_{1,2}(\rho,s)$ and $\zeta_{*}\equiv\zeta_{*}(\rho)$.
Figure 4. Schematic plot of $F(k;\rho)$ along the imaginary axis. The function is convex with the minimum at $\zeta_{*}\equiv\zeta_{*}(\rho)$ which is negative if $\mu_{1}(\rho)<0$ and positive if $\mu_{1}(\rho)>0$. If $s\,c(\rho)<1$, then there are two real solutions $\zeta_{1,2}\equiv\zeta_{1,2}(\rho,s)$ to (39) located on opposite sides of the origin.
The location of $\zeta_{1,2}$ defines the strip in the $k$-plane, where the logarithmic term in the integrand (9) is analytic. Indeed, due to (35) we have
$\left|s\,c(\rho)\,F(k;\rho)\right|<1,\quad\,\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2}).$ (40)
Note that (40) implies that there are two branching points of the logarithm at $k=\mathrm{i}\zeta_{1,2}$ (see Fig. 5). In principle, there may be other singularities in the logarithmic term arising from the non-real solutions of (39), but they all lie outside the horizontal strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$. With this we conclude the list of properties of $F(k;\rho)$ that are important for the analysis to come.
Figure 5. The structure of $\log[1-sc(\rho)F(k;\rho)]$ in the complex $k$-plane. The function is analytic in the horizontal strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$ with branching points at $k=\mathrm{i}\zeta_{1,2}$. The edges of the strip $\zeta_{1,2}$ are two real solutions of (39) as shown in Fig. 4. Note that (39) may have other non-real solutions, leading to more singularities in the logarithmic term, but due to (40) these singularities lie in the shaded region.
3.2. A toy example
Before performing an analytic continuation of $\phi^{\pm}(\lambda;\rho,s)$ to the $\lambda$-complex plane, let us consider a simplified version of the problem. Specifically, we define two auxiliary functions
$\psi^{\pm}(\lambda)=\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\,\frac{1}{\lambda\pm\mathrm{i}k}\log\left[(k-\mathrm{i}A)(k+\mathrm{i}B)\right]\right],$ (41)
where $A$ and $B$ are positive real parameters and the integral is understood in terms of the principal value.
The functions $\psi^{\pm}(\lambda)$ resemble $\phi^{\pm}(\lambda;\rho,s)$ in structure, but are sufficiently simple to allow explicit evaluation for all complex $\lambda$. They thus offer an illustrative example in which to expose the difference between a naive extension of (41) and a proper analytic continuation. Moreover, there is a pragmatic motive for this brief detour. Similar integrals will reappear in the analysis of first-passage properties, and it is prudent to compute them here, killing two birds with one well-aimed stone.
The standard method for evaluating the integrals of the form (41) involves extending them into the complex $k$-plane and closing the contour either in the lower or in the upper half-plane. If the contribution from the semicircle vanishes at infinity, the integral is determined by the singularities inside the contour. However, in the form (41) it is not apparent why the contributions from the arcs decay sufficiently fast (faster than $\left|k\right|^{-1}$). For this reason, we modify this representation so that the standard procedure applies.
When integrating (41) over the real line, the odd part of the integrand cancels due to symmetry, and only the even part remains. Specifically, we can write
$\psi^{\pm}(\lambda)=\exp\left[-\frac{1}{4\pi}\int_{-\infty}^{\infty}\mathrm{d}k\left\{\frac{\log\left[(k-\mathrm{i}A)(k+\mathrm{i}B)\right]}{\lambda\pm\mathrm{i}k}+\frac{\log\left[(k+\mathrm{i}A)(k-\mathrm{i}B)\right]}{\lambda\mp\mathrm{i}k}\right\}\right].$ (42)
Rearranging the terms in the integrand we obtain
$\log\left[\psi^{\pm}(\lambda)\right]=-\frac{1}{4\pi}\int_{-\infty}^{\infty}\mathrm{d}k\frac{\lambda}{\lambda^{2}+k^{2}}\log\left[\left(k^{2}+A^{2}\right)\left(k^{2}+B^{2}\right)\right]\\ \mp\frac{1}{4\pi}\int_{-\infty}^{\infty}\mathrm{d}k\frac{\mathrm{i}k}{\lambda^{2}+k^{2}}\log\left[\frac{(k+\mathrm{i}A)(k-\mathrm{i}B)}{(k-\mathrm{i}A)(k+\mathrm{i}B)}\right].$ (43)
Finally, integrating the second term by parts yields a decomposition of $\log\psi^{\pm}(\lambda)$ into four integrals of essentially identical form:
$\log\left[\psi^{\pm}(\lambda)\right]=-\frac{1}{4\pi}\int_{-\infty}^{\infty}\mathrm{d}k\left(\frac{\lambda}{\lambda^{2}+k^{2}}\log\left[k^{2}+A^{2}\right]+\frac{\lambda}{\lambda^{2}+k^{2}}\log\left[k^{2}+B^{2}\right]\right)\\ \pm\frac{1}{4\pi}\int_{-\infty}^{\infty}\mathrm{d}k\left(\frac{A}{A^{2}+k^{2}}\log\left[k^{2}+\lambda^{2}\right]-\frac{B}{B^{2}+k^{2}}\log\left[k^{2}+\lambda^{2}\right]\right).$ (44)
Each integral in (44) can now be evaluated using standard complex analysis techniques, by closing the contour in the complex $k$-plane. In particular, all integrals reduce to the application of the identity
$\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{b}{b^{2}+k^{2}}\log(k^{2}+a^{2})\mathrm{d}k=\frac{b}{\sqrt{b^{2}}}\log\frac{b+a}{b-a}+\frac{a}{\sqrt{a^{2}}}\log\left[b^{2}-a^{2}\right],$ (45)
where $a$ and $b$ are arbitrary complex parameters, and the integral is taken as a principal value.
Applying (45) to the integrals in (44) after algebraic manipulations, we find that
[…]
From (46) it is evident that the analytic continuations $\psi_{\text{a.c.}}^{\pm}(\lambda)$ of the functions $\psi^{\pm}(\lambda)$ from the positive half of the real line $\lambda\in(0;\infty)$ to the full $\lambda$-plane are given by:
$\psi^{+}_{\text{a.c.}}(\lambda)=\frac{1}{B+\lambda},\qquad\psi^{-}_{\text{a.c.}}(\lambda)=\frac{1}{A+\lambda},\qquad\lambda\in\mathbb{C}.$ (47)
Not only do the functions $\psi^{\pm}(\lambda)$, computed in (46) via the original integral expression (41) for arbitrary $\lambda$, differ from their analytic continuations (47), they are in fact discontinuous at $\mathrm{Re}(\lambda)=0$. Having this explicit computation in mind, we now show how to perform the analytic continuation at the level of the integral expression (41).
We first focus on the function $\psi^{+}(\lambda)$. The analytic structure of the integrand for $\psi^{+}(\lambda)$ in the $k$-plane is relatively simple: it has a branch cut from the logarithm and a pole at $k=\mathrm{i}\lambda$. When $\mathrm{Re}(\lambda)\to 0^{+}$, this pole approaches the real axis from above, falling on the integration contour when $\mathrm{Re}(\lambda)=0$ and crossing it to the lower half-plane for $\mathrm{Re}(\lambda)<0$. This is exactly the mechanism responsible for the discontinuity of $\psi^{+}(\lambda)$.
Figure 6. Structure of the integrand for $\psi^{+}(\lambda)$ in (41) and the deformed contour $\mathcal{C}$ used for the analytic continuation. For $\mathrm{Re}(\lambda)>0$, the contour coincides with the real line. For $\mathrm{Re}(\lambda)=0$, the contour circumvents the pole as a semicircle. For $\mathrm{Re}(\lambda)<0$, the contour encircles the pole in the lower half-plane.
In order to analytically continue $\psi^{+}(\lambda)$, we deform the contour of integration to circumvent the pole as shown in Fig. 6 and compute the integral explicitly. The analytic continuation of the other function $\psi^{-}(\lambda)$ is performed in essentially the same way. The only difference is that the pole is located at $k=-\mathrm{i}\lambda$, and when $\mathrm{Re}(\lambda)$ decreases from positive to negative, it crosses the real line in the opposite direction.
Let us now compute the integral over the deformed contours for $\mathrm{Re}(\lambda)\leq 0$. When $\mathrm{Re}(\lambda)=0$, the pole lies precisely on the integration contour and the principal value integral differs from the analytic continuation by half the residue at $k=\pm\mathrm{i}\lambda$,
$\mathrm{Re}(\lambda)=0:\quad\log\left[\psi_{\text{a.c.}}^{\pm}(\lambda)\right]=\log\left[\psi^{\pm}(\lambda)\right]\\ \mp\pi\mathrm{i}\operatorname*{Res}_{k=\pm\mathrm{i}\lambda}\left[\frac{1}{2\pi}\frac{\log\left[(k-\mathrm{i}A)(k+\mathrm{i}B)\right]}{\lambda\pm\mathrm{i}k}\right].$ (48)
For $\mathrm{Re}(\lambda)<0$, the contour consists of the real line, two vertical segments, and a circle around the pole at $k=\pm\mathrm{i}\lambda$. The integrals over the vertical segments cancel one another, and the circle integral is equal to the residue at $k=\pm\mathrm{i}\lambda$. Thus, the analytic continuation differs from the real-line integral by the full residue, namely,
$\mathrm{Re}(\lambda)<0:\quad\log\left[\psi_{\text{a.c.}}^{\pm}(\lambda)\right]=\log\left[\psi^{\pm}(\lambda)\right]\\ \mp 2\pi\mathrm{i}\operatorname*{Res}_{k=\pm\mathrm{i}\lambda}\left[\frac{1}{2\pi}\frac{\log\left[(k-\mathrm{i}A)(k+\mathrm{i}B)\right]}{\lambda\pm\mathrm{i}k}\right].$ (49)
Computing the residues explicitly, we obtain the analytic continuations $\psi_{\text{a.c.}}^{\pm}(\lambda)$ of $\psi^{\pm}(\lambda)$,
[…]
Comparing the analytic continuation (50), derived directly from the integral representation (41), with the continuation (47) obtained after evaluating the integrals in (46), we see that the two match exactly. This confirms the validity of the contour deformation procedure and the consistency of the analytic continuation.
An important remark here is that the above procedure implicitly assumes that the deformed contour $\mathcal{C}$ does not cross the branch cut of the logarithm. This assumption is however easily lifted. Consider a strip $\mathrm{Im}(k)\in(-a,a)$, where $a=\min(A,B)$. Within this strip, the logarithmic term in the integrand (41) is analytic. If $\left|\,\mathrm{Re}(\lambda)\right|<a$, then the deformed contour lies within $\mathrm{Im}(k)\in(-a,a)$ where there are no singularities. Hence the analytic continuation (50) is valid in this strip. To extend the continuation to the full region $\mathrm{Re}(\lambda)<0$, we combine (50) with (41) to obtain the functional relation
$\mathrm{Re}(\lambda)\in(0,a):\quad\psi^{\mp}_{\text{a.c.}}(\lambda)\psi^{\pm}_{\text{a.c.}}(-\lambda)=\frac{\psi^{\mp}(\lambda)\psi^{\pm}(-\lambda)}{(A\mp\lambda)(B\pm\lambda)}=\frac{1}{(A\mp\lambda)(B\pm\lambda)}.$ (51)
Now, using (51) as the definition of the analytic continuation for all $\lambda$, we obtain
[…]
Clearly (52) and (50) coincide.
3.3. Continuation in the $\lambda$-plane
Now we proceed to the analytic continuation of $\phi^{\pm}(\lambda;\rho,s)$ to the $\lambda$-plane, assuming that $s\in(0,1)$ and $\rho>0$. Recall that according to (9) the functions $\phi^{\pm}(\lambda;\rho,s)$ are given by
$\phi^{\pm}(\lambda;\rho,s)=\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\;\frac{1}{\lambda\pm\mathrm{i}k}\log\left[1-s\,c(\rho)F(k;\rho)\right]\right].$ (53)
The similarity with the example (41) is apparent. There is only one $\lambda$-dependent singularity, i.e., a pole at $k=\pm\mathrm{i}\lambda$. The analytic structure of the logarithm on the other hand depends on the exact form of the function $F(k;\rho)$ and can be rather involved. Fortunately, there is no need to analyze it in detail, and we can focus on the strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$.
The analytic continuation of $\phi^{\pm}(\lambda;\rho,s)$ is performed in three steps. The first step is to continue $\phi^{\pm}(\lambda;\rho,s)$ to the region $\mathrm{Re}(\lambda)>0$ by simply extending the definition (53). The second step is to consider a strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$ where the logarithmic term has no singularities (see Fig. 5). Within this strip we can freely deform the contour of integration to encircle the pole in the same way as was done in the previous example (see Fig. 6). Then, separating the contribution from the pole and computing the residue we obtain the functional relation
$\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)\;\phi_{\text{a.c.}}^{\mp}(-\lambda;\rho,s)=\frac{1}{1-s\,c(\rho)\,F(\pm\mathrm{i}\lambda;\rho)}.$ (54)
The third and final step is to use (54) as the definition of $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ for all $\mathrm{Re}(\lambda)<0$. This procedure results in the analytic continuation
[…]
Recall that $F(k;\rho)$ is, in general, defined for $\mathrm{Im}(k)\in(\zeta_{-},\zeta_{+})$ hence the lower bound for $\mathrm{Re}(\lambda)$ in (55).
The functions $\phi^{\pm}(\lambda;\rho,s)$ are regular when $\mathrm{Re}(\lambda)\neq 0$. Consequently their analytic continuations have no singularities in the right half-plane $\mathrm{Re}(\lambda)>0$ and all singularities that may appear in the left half-plane $\mathrm{Re}(\lambda)<0$ are those of the prefactor. As for the imaginary line $\mathrm{Re}(\lambda)=0$, the analytic continuations are regular there by construction.
Figure 7. Analytic structure of $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ in the $\lambda$-plane. Left panel: (55) defines the function $\phi^{+}_{\text{a.c.}}(\lambda;\rho,s)$ in the region $\mathrm{Re}(\lambda)>\zeta_{-}$; the rightmost singularity is a pole at $\lambda=\zeta_{1}(\rho,s)$. Right panel: (55) defines the function $\phi^{-}_{\text{a.c.}}(\lambda;\rho,s)$ in the region $\mathrm{Re}(\lambda)>-\zeta_{+}$; the rightmost singularity is a pole at $\lambda=-\zeta_{2}(\rho,s)$.
In other words, the functions $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ have singularities at $\lambda^{\pm}_{*}$ in the left half-plane, which are solutions of
$1-s\,c(\rho)\,F(\pm\mathrm{i}\lambda^{\pm}_{*};\rho)=0,\quad\,\mathrm{Re}(\lambda^{\pm}_{*})<0,$ (56)
This is the same equation (up to a sign of the argument) as (39). Recall that for the “++” sign, it admits two real solutions $\zeta_{1,2}(\rho,s)$ with $\zeta_{1}<0<\zeta_{2}$ and, due to (40), has no solutions in the strip $\mathrm{Re}(\lambda)\in(\zeta_{1},\zeta_{2})$. Therefore, the rightmost singularity of $\phi^{+}_{\text{a.c.}}(\lambda;\rho,s)$ in the left half-plane is a pole at $\lambda=\zeta_{1}(\rho,s)$. Similarly, the rightmost singularity of $\phi^{-}_{\text{a.c.}}(\lambda;\rho,s)$ is a pole at $\lambda=-\zeta_{2}(\rho,s)$ (see Fig. 7).
4. Survival probability $S_{\infty}(X_{0})$
Having performed the continuation of $\phi^{\pm}(\lambda;\rho,s)$ into $\lambda$-plane, the next steps are to consider the continuations into $s$- and $\rho$-planes, as the analytic structure in these planes provides us with the asymptotic behaviors of $S_{N}(n\,|\,X_{0})$ and $S_{T}(\tau\,|\,X_{0})$. It is, however, more instructive to start with a more robust quantity: the probability to survive indefinitely $S_{\infty}(X_{0})$. The Laplace transform $\hat{S}_{\infty}(\lambda)$ of $S_{\infty}(X_{0})$ is given by
$\hat{S}_{\infty}(\lambda)=\int_{0}^{\infty}\mathrm{d}X_{0}e^{-\lambda X_{0}}\left(1-\int_{0}^{\infty}\mathrm{d}\tau\sum_{n=0}^{\infty}\mathbb{P}[\tau,n\,|X_{0}]\right)=\frac{1}{\lambda}-\lim_{\begin{subarray}{c}\rho\to 0\\ s\to 1\end{subarray}}\hat{Q}(\lambda;\rho,s).$ (57)
Finding the exact expression for $\hat{S}_{\infty}(\lambda)$ requires only computing the limit $s\to 1$, $\rho\to 0$ of $\hat{Q}(\lambda;\rho,s)$ rather than performing the full analytic continuation to the $\rho$- and $s$-planes.
In this section we compute the limit in (57) and show that there are indeed two different regimes separated by the critical drift $\alpha_{c}$: the absorption regime, where $S_{\infty}(X_{0})=0$ and the survival regime, where $S_{\infty}(X_{0})>0$. For the survival regime we compute the first two terms of the asymptotic expansion of $S_{\infty}(X_{0})$ when $X_{0}\to 0$ and the rate of exponential decay for $X_{0}\to\infty$.
4.1. Explicit expression for the Laplace transform
Recall that the analytic structure of integrals entering $\phi_{\text{a.c.}}^{\pm}(\lambda;\rho,s)$ in the $k$-plane described in Sec. 3 (see Fig. 5) is valid for $s<1$ and $\rho>0$. In the limit $s\to 1$ and $\rho\to 0$ the structure changes, as one of the branching points of the logarithm falls onto the real line inducing the singularity at $k=0$. Specifically, if $\langle M-\alpha t\rangle<0$ then $\zeta_{2}=0$, and if $\left\langle M-\alpha t\right\rangle>0$ then $\zeta_{1}=0$ (see Fig. 4). At the same time, as $\lambda\to 0$ the pole $k=-\mathrm{i}\lambda$ in $\phi^{+}_{\text{a.c.}}(\lambda;\rho,s)$ also falls onto the real line at $k=0$. Recall that $\phi_{\text{a.c.}}^{+}(0;\rho,s)$ is one of the factors in the representation (8) for $\hat{Q}(\rho,s\,|\,\lambda)$, hence, in the limit $\rho\to 0$ and $s\to 1$, the coincidence of two singularities at the origin may induce a non-trivial behavior and needs to be dealt with.
Since the problem may arise due to the singularity at $k=0$, we preemptively avoid potential complications by explicitly factoring out the possibly singular behavior from the integral before taking the limit. To simplify the notation we will focus on the case $\lambda>0$. The continuation to $\lambda<0$ can be found straightforwardly from (55). First, we formally rewrite (53) as:
$\phi^{\pm}(\lambda;\rho,s)=\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\;\frac{1}{\lambda\pm\mathrm{i}k}\log\left[\frac{1-s\,c(\rho)F(k;\rho)}{1-s\,c(\rho)F_{2}(k;\rho)}\right]\right]\\ \times\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}k}{\lambda\pm\mathrm{i}k}\log\left[1-s\,c(\rho)F_{2}(k;\rho)\right]\right],$ (58)
where
$F_{2}(k;\rho)=1+\mathrm{i}\,\mu_{1}(\rho)\;k-\frac{1}{2}\mu_{2}(\rho)k^{2}.$ (59)
Note that (58) amounts to adding and subtracting the term corresponding to the expansion of the integrand in the vicinity of $k=0$. This term can easily be computed since
$1-s\,c(\rho)\,F_{2}(k;\rho)=\frac{1}{2}s\,c(\rho)\,\mu_{2}(\rho)\big(k-\mathrm{i}k_{+}^{*}(\rho,s)\big)\big(k+\mathrm{i}k_{-}^{*}(\rho,s)\big),$ (60)
with
$k_{\pm}^{*}(\rho,s)=\pm\frac{\mu_{1}(\rho)}{\mu_{2}(\rho)}+\sqrt{\left(\frac{\mu_{1}(\rho)}{\mu_{2}(\rho)}\right)^{2}+\frac{2}{\mu_{2}(\rho)}\left(\frac{1}{s\,c(\rho)}-1\right)}.$ (61)
Applying (60) to the second term in (58) yields an integral that, up to a constant, has the same form as the auxiliary function $\psi^{+}(\lambda)$ given by (41), with $A$ replaced by $k_{+}^{*}$ and $B$ replaced by $k_{-}^{*}$. Using (46) we obtain for $\mathrm{Re}(\lambda)>0$
$\mathrm{Re}(\lambda)>0:\quad\phi^{\pm}(\lambda;\rho,s)=\frac{1}{\lambda+k_{\mp}^{*}(\rho,s)}\sqrt{\frac{2}{s\,c(\rho)\,\mu_{2}(\rho)}}\\ \times\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\;\frac{1}{\lambda\pm\mathrm{i}k}\log\left[\frac{1-s\,c(\rho)\,F(k;\rho)}{1-s\,c(\rho)\,F_{2}(k;\rho)}\right]\right].$ (62)
Although this representation is obtained for $\mathrm{Re}(\lambda)>0$ it can be readily extended to the full $\lambda$-plane. Such an extension consists of combining the prefactors already found in (50) and (55). The resulting expressions are rather cumbersome and hence we do not present them here explicitly, except for one observation. Specifically, since there is no pole at $k=0$ in the integrand when $\lambda=0$, the analytic continuation $\phi_{\text{a.c.}}^{\pm}(\lambda;\rho,s)$ evaluated at $\lambda=0$ is given by
$\phi^{\pm}_{\text{a.c.}}(0;\rho,s)=\frac{1}{k_{\mp}^{*}(\rho,s)}\sqrt{\frac{2}{s\,c(\rho)\,\mu_{2}(\rho)}}\\ \times\exp\left[\mp\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\;\frac{1}{\mathrm{i}k}\log\left[\frac{1-s\,c(\rho)\,F(k;\rho)}{1-s\,c(\rho)\,F_{2}(k;\rho)}\right]\right].$ (63)
The main advantage of (63) is that the integral is regular as $s\to 1$, $\rho\to 0$ and the nontrivial dependence is encoded solely in the prefactor.
Substituting now (62) and (63) into (8) we immediately obtain that the Laplace transform $\hat{Q}(\rho,s\,|\,\lambda)$ can be written as
$\mathrm{Re}(\lambda)\geq 0:\quad\hat{Q}(\rho,s\,|\,\lambda)=\frac{1}{\lambda+\frac{\rho}{\alpha}}\\ -\frac{1-s\,c(\rho)}{\lambda+\frac{\rho}{\alpha}}\frac{2}{s\,c(\rho)\,\mu_{2}(\rho)}\frac{1}{k_{-}^{*}(\rho,s)\left(\lambda+k_{+}^{*}(\rho,s)\right)}e^{I(\lambda;\rho,s)},$ (64)
where
$I(\lambda;\rho,s)=-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\,\left(\frac{1}{\mathrm{i}k}+\frac{1}{\lambda+\frac{\rho}{\alpha}-\mathrm{i}k}\right)\log\frac{1-s\,c(\rho)\,F(k;\rho)}{1-s\,c(\rho)\,F_{2}(k;\rho)}.$ (65)
The representation (64) is convenient to compute the limit in (57) and thereby the Laplace transform $\hat{S}_{\infty}(\lambda)$.
In the limit $s\to 1$, $\rho\to 0$ the integral (65) is regular and the nontrivial behavior is contained in the prefactor. A simple calculation shows that
$\lim_{\rho\to 0,s\to 1}\left[\frac{1-s\,c(\rho)}{k_{-}^{*}(\rho,s)}\frac{1}{\lambda+k_{+}^{*}(\rho,s)}\frac{2}{s\,c(\rho)\,\mu_{2}(\rho)}\right]=\begin{cases}0,&\mu_{1}(0)\leq 0,\\ \frac{1}{1+\lambda\frac{\mu_{2}(0)}{2\mu_{1}(0)}},&\mu_{1}(0)>0.\end{cases}$ (66)
Hence, recalling the definition (30) of the moments $\mu_{\ell}(\rho)$, we obtain
$\left\langle M-\alpha t\right\rangle\leq 0:$ $\qquad\hat{S}_{\infty}(\lambda)=0,$ (67)
$\left\langle M-\alpha t\right\rangle>0:$ $\qquad\hat{S}_{\infty}(\lambda)=\frac{1}{\lambda}\frac{1}{1+\lambda\frac{\left\langle(M-\alpha t)^{2}\right\rangle}{2\left\langle M-\alpha t\right\rangle}}e^{I(\lambda;0,1)}.$ (68)
From (67) and (68) the existence of survival and absorption regimes is evident, and so is the critical value of the drift (6) separating these two regimes, i.e., $\alpha_{c}=\langle M\rangle/\langle t\rangle$.
The survival probability at the infinite time $S_{\infty}(X_{0})$ in the survival regime can in principle be obtained by inverting the Laplace transform
$\mathrm{Re}(\lambda)\geq 0:\qquad\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}S_{\infty}(X_{0})=\frac{1}{\lambda}\frac{1}{1+\lambda\frac{\left\langle(M-\alpha t)^{2}\right\rangle}{2\left\langle M-\alpha t\right\rangle}}e^{I(\lambda;0,1)}.$ (69)
The inversion is, however, unfeasible unless the probability distributions $p(t)$ and $q(M)$ have a specific form. Instead, we extract the asymptotic behavior of $S_{\infty}(X_{0})$ for $X_{0}\to\infty$ and $X_{0}\to 0$.
4.2. Asymptotic behavior
The behavior as $X_{0}\to 0$ is governed by the $\lambda\to\infty$ expansion of $\hat{S}_{\infty}(X_{0})$. The detailed calculation is provided in Appendix A, and here we present only the result. Specifically, substituting the expansion $I(\lambda;0,1)$ given by (207) into (69), after algebraic manipulations, yields
$\hat{S}_{\infty}(\lambda)\underset{\lambda\to\infty}{=}\frac{1}{\lambda}\sqrt{\frac{2\langle M-\alpha t\rangle^{2}}{\left\langle(M-\alpha t)^{2}\right\rangle}}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\;\log\left|1-F(k;0)\right|\right)+O\left(\frac{1}{\lambda^{2}}\right),$ (70)
which, according to Tauberian theorems, translates into
$S_{\infty}(X_{0})\underset{X_{0}\to 0}{\sim}\sqrt{\frac{2\langle M-\alpha t\rangle^{2}}{\langle(M-\alpha t)^{2}\rangle}}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{X_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\log\left|1-F(k;0)\right|\right).$ (71)
As a consistency check, note that if $\alpha=\alpha_{c}$, then $\langle M-\alpha_{c}t\rangle=0$ and the prefactor in (71) is zero, which agrees with $S_{\infty}(X_{0})=0$ anticipated in (67).
The behavior of $S_{\infty}(X_{0})$ as $X_{0}\to\infty$ can be inferred from the singularity structure of $\hat{S}_{\infty}(\lambda)$ in the $\lambda$-plane. The representation (69) is regular for $\mathrm{Re}(\lambda)>0$ thus, this singularity lies in $\mathrm{Re}(\lambda)\leq 0$. By either keeping track of the analytic structure in (62) and combining the prefactors found in (50) and (55) or, alternatively, by repeating the same arguments as presented in Sec. 3.3 at the level of the representation (68) we find that
$\mathrm{Re}(\lambda)\leq 0:\qquad\hat{S}_{\infty}(\lambda)=\frac{1-F_{2}(-\mathrm{i}\lambda;0)}{1-F(-\mathrm{i}\lambda;0)}\frac{1}{\lambda}\frac{1}{1+\lambda\frac{\left\langle(M-\alpha t)^{2}\right\rangle}{2\left\langle M-\alpha t\right\rangle}}e^{I(\lambda;0,1)},$ (72)
or, using the explicit form of $F_{2}(k;\rho)$ as in (59),
$\mathrm{Re}(\lambda)\leq 0:\qquad\hat{S}_{\infty}(\lambda)=-\frac{\left\langle M-\alpha t\right\rangle}{1-F(-\mathrm{i}\lambda;0)}e^{I(\lambda;0,1)}.$ (73)
As we have argued before in Sec. 3 the denominator has two real solutions at $\lambda=-\zeta_{1,2}$ (see right panel of Fig. 4) and no solutions in between these two values. In the limit $\rho\to 0$ and $s\to 1$ we have $\zeta_{1}=0$, hence in (73) there are two poles: one at $\lambda=0$, and the other at $\lambda=-\zeta_{2}$.
The leading behavior of $S_{\infty}(X_{0})$ as $X_{0}\to\infty$ is governed by the pole at $\lambda=0$. Expanding (73) as $\lambda\to 0$ we find that
$\hat{S}_{\infty}(\lambda)=\frac{1}{\lambda}+O\left(1\right),\qquad\lambda\to 0,$ (74)
which gives the asymptotic value of $S_{\infty}(X_{0})$
$\lim_{X_{0}\to\infty}S_{\infty}(X_{0})=1.$ (75)
The result (75) is indeed very natural. In the survival regime $\langle M-\alpha t\rangle>0$ so that on average the process moves away from the origin. Therefore, the further away from the origin it starts, the lower is the probability of the first-passage event. In the limit $X_{0}\to\infty$ this probability tends to $1$. The more interesting question is how this asymptotic value is approached.
The subleading corrections can be inferred from the singularities of $\hat{S}_{\infty}(\lambda)$. The first of them corresponds to the singularity with the largest real part. In the considered case, this is the pole at $\lambda=-\zeta_{2}(0,1)$. Introducing the notation $R\equiv\zeta_{2}(0,1)$ we write
$S_{\infty}(X_{0})-1\underset{X_{0}\to\infty}{\asymp}e^{-RX_{0}}.$ (76)
Recall that $R$ is the smallest positive solution of $F(\mathrm{i}R;0)=1$.
As a final remark here, we should mention that the behavior (76) is actually well-known in the financial literature, in which $S_{\infty}(X_{0})$ denotes the probability of the ultimate ruin. In this context, the condition $F(\mathrm{i}R;0)=1$ is essentially equivalent to what is known as the Lundberg fundamental equation [5]. Here we obtain this decay rate as a byproduct, which both serves as a consistency check and demonstrates the power of the representation (8).
Now we proceed with the analytic continuation in the $s$- and $\rho$-planes. For simplicity and to align with the goal of studying first-passage properties, we will be performing the continuations directly for the Laplace transforms of the survival probabilities $S_{N}(n\,|\,X_{0})$ and $S_{T}(\tau\,|\,X_{0})$, rather than for the functions $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$.
5. Survival probability $S_{N}(n\,|\,X_{0})$
Let us first consider the probability of the process staying positive up to the $n$-th jump. The Laplace transform of this probability
$\hat{S}_{N}(s\,|\,\lambda)\equiv\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\sum_{n=0}^{\infty}s^{n}S_{N}(n\,|\,X_{0})$ (77)
can be represented in terms of $\hat{Q}(\rho,s\,|\,\lambda)$ as
$\hat{S}_{N}(s\,|\,\lambda)=\left.\frac{1}{1-s}\left(\frac{1}{\lambda}-\hat{Q}(\rho,s\,|\,\lambda)\right)\right|_{\rho=0}.$ (78)
Equivalently, using (8) and replacing $\phi^{\pm}(\lambda;\rho,s)$ with their respective analytic continuations (55), we find
$\hat{S}_{N}(s\,|\,\lambda)=\frac{1}{\lambda}\phi^{+}_{\text{a.c.}}(0;\rho,s)\;\phi^{-}_{\text{a.c.}}\left(\lambda+\frac{\rho}{\alpha};\rho,s\right)\Big|_{\rho=0}.$ (79)
In contrast to Sec. 4 it is more convenient to use explicit regularization for $\phi^{+}_{\text{a.c.}}(0;\rho,s)$ in (53), rather than utilizing (62) or treating the integral in (53) in the principal value sense. Specifically, we will use the following representation:
$\hat{S}_{N}(s\,|\,\lambda)=\lim_{\epsilon\to 0}\frac{1}{\lambda}\phi^{+}_{\text{a.c.}}(\epsilon;\rho,s)\;\phi^{-}_{\text{a.c.}}\left(\lambda+\frac{\rho}{\alpha};\rho,s\right)\Big|_{\rho=0}.$ (80)
The analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ in the $\lambda$-plane follows directly from those of $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ as given in (55). In this section we focus on its behavior in the $s$-plane.
As will be shown shortly, the analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ depends on the sign of $\mu_{1}(0)=\langle M-\alpha t\rangle$, since $\mu_{1}(0)>0$ and $\mu_{1}(0)<0$ correspond to the survival and absorption regimes respectively, and $\mu_{1}(0)=0$ is a critical point. We will study these cases separately.
5.1. Survival regime
We start with the survival regime, i.e., the case where $\mu_{1}(0)>0$. In this scenario, we expect a finite probability that the process stays positive indefinitely and the first-passage never occurs. By analyzing the analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ in the $s$-plane we show that $S_{\infty}(X_{0})>0$ and that $S_{N}(n\,|\,X_{0})$ approaches $S_{\infty}(X_{0})$ exponentially fast with the rate given by $\xi_{N}(\alpha)$ as in (14).
Analytic continuation in the $s$-plane
Assume for the moment that $\lambda>0$ (the case $\lambda\leq 0$ will be considered shortly). Then according to (55) we have:
$\lambda>0:\quad\phi_{\text{a.c.}}^{\pm}(\lambda;0,s)=\phi^{\pm}(\lambda;0,s).$ (81)
Hence (80) implies
$\hat{S}_{N}(s\,|\,\lambda)=\lim_{\epsilon\to 0}\frac{1}{\lambda}\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\,\left(\frac{1}{\epsilon+\mathrm{i}k}+\frac{1}{\lambda-\mathrm{i}k}\right)\log\left[1-sF(k;0)\right]\right].$ (82)
The continuation in the $s$-plane is performed by deforming the contour of integration in the $k$-plane. The analytic structure of the integrand is as follows (see Fig. 8 for the schematic illustration): The logarithm is analytic within the strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$ and has branch cuts at $k=\mathrm{i}\zeta_{1,2}$, where $\zeta_{1,2}\equiv\zeta_{1,2}(0,s)$ are the real solutions of (39) with $\rho=0$ (see Fig. 4 and Fig. 5). The term in parentheses contributes two simple poles at $k=\mathrm{i}\epsilon$ and $k=-\mathrm{i}\lambda$.
Figure 8. Schematic representation of the analytic structure of the integrand in (82) in the complex $k$-plane within the survival regime. The integrand features two branch points at $k=\mathrm{i}\zeta_{1,2}$ and two poles at $k=\mathrm{i}\epsilon$ and $k=-\mathrm{i}\lambda$. Upon continuation to $s>1$, the branch points pinch the shifted contour $\mathcal{C}_{*}$ inducing a singularity at $s=s_{*}$ defined by (85) in the $s$-plane.
Note that $c(0)=1$, and thus as $s\to 1$, we have $\zeta_{1}(0,s)\to 0$. This implies that the branch point at $\mathrm{i}\zeta_{1}(0,s)$ moves onto the contour of integration at $k=0$. Simultaneously, the pole at $k=\mathrm{i}\epsilon$ also approaches the contour at $k=0$ as $\epsilon\to 0$. This is exactly the coincidence of singularities studied in Sec. 4.
To perform the continuation to the region $s>1$, we shift the contour of integration in (82) from the real line to a horizontal contour $\mathcal{C}_{*}$ such that $\mathrm{Im}(k)=\zeta_{*}(0)$, where $\zeta_{*}(\rho)$ is the location of the minimum of $F(k;\rho)$ along the imaginary axis (see Fig. 4). Recall that if $\mu_{1}(0)>0$, then $\zeta_{*}(0)>0$, and when shifting the contour we must account for the residue at $k=\mathrm{i}\epsilon$. The resulting expression is
$\hat{S}_{N}(s\,|\,\lambda)=\lim_{\epsilon\to 0}\Biggl\{\frac{1}{\lambda}\frac{1}{1-sF(\mathrm{i}\epsilon;0)}\\ \times\exp\left[-\frac{1}{2\pi}\int_{\mathcal{C}_{*}}\mathrm{d}k\,\left(\frac{1}{\epsilon+\mathrm{i}k}+\frac{1}{\lambda-\mathrm{i}k}\right)\log\left[1-sF(k;0)\right]\right]\Biggr\}.$ (83)
Taking the limit $\epsilon\to 0$ is now straightforward and we have
$\hat{S}_{N}(s\,|\,\lambda)=\frac{1}{\lambda}\frac{1}{1-s}\exp\left[-\frac{1}{2\pi}\int_{\mathcal{C}_{*}}\mathrm{d}k\,\left(\frac{1}{\mathrm{i}k}+\frac{1}{\lambda-\mathrm{i}k}\right)\log\left[1-sF(k;0)\right]\right].$ (84)
As the value of $s$ increases, the branch points at $\mathrm{i}\zeta_{1,2}$ approach the contour $\mathcal{C}_{*}$, finally pinching once $s$ reaches a critical value $s_{*}$. The pinch occurs when
$s_{*}:\quad\zeta_{1}(0;s_{*})=\zeta_{2}(0;s_{*})=\zeta_{*}(0),$ (85)
or, equivalently, when (39) has one real solution, i.e.,
$\frac{1}{s_{*}}=F\left(\mathrm{i}\zeta_{*}(0);0\right)=\min_{k_{I}}F\left(\mathrm{i}k_{I};0\right).$ (86)
The analytic structure of (84) in the $s$-plane is now clear. The prefactor gives rise to a simple pole at $s=1$. The integral remains analytic for $\left|s\right|<s_{*}$, and has a singularity at $s=s_{*}$, which a more careful analysis reveals to be a branch point, though its precise nature is not essential here. The expression (84) therefore provides an analytic continuation valid for $\left|s\right|<s_{*}$. The only step remaining is to extend the above derivation to the case $\mathrm{Re}(\lambda)\leq 0$, in which according to (55), we have
$\lambda<0:\quad\phi_{\text{a.c.}}^{\pm}(\lambda;0,s)=\frac{1}{1-sF(\pm\mathrm{i}\lambda;0)}\phi^{\pm}(\lambda;0,s).$ (87)
The analysis proceeds the same way. The only subtlety here is that the pole at $k=-\mathrm{i}\lambda$ lies in the upper half-plane. If $\lambda\in\left(-\zeta_{*},0\right)$, then while shifting the contour from the real line to $\mathcal{C}_{*}$ we must account for two poles: at $k=\mathrm{i}\epsilon$ and at $k=-\mathrm{i}\lambda$, with the contribution from the latter canceling out the prefactor.
Combining all the arguments above we arrive at the analytic continuation of $\hat{S}_{N}(s\,|\,\lambda)$ valid within the disk of radius $s_{*}$ defined by (86), namely,
$\alpha<\alpha_{c}:\qquad\hat{S}_{N}(s\,|\,\lambda)=\left\{\begin{aligned} &\frac{1}{\lambda}\frac{\mathcal{G}_{N}(s\,|\,\lambda)}{1-s},\quad\lambda>-\zeta_{*}(0),\\ &\frac{1}{1-sF\left(-\mathrm{i}\lambda;0\right)}\frac{1}{\lambda}\frac{\mathcal{G}_{N}(s\,|\,\lambda)}{1-s},\quad\lambda<-\zeta_{*}(0),\end{aligned}\right$. (88)
where we used a notation
$\mathcal{G}_{N}(s\,|\,\lambda)=\exp\left[-\frac{1}{2\pi}\int_{\mathcal{C}_{*}}\mathrm{d}k\,\left(\frac{1}{\mathrm{i}k}+\frac{1}{\lambda-\mathrm{i}k}\right)\log\left[1-sF(k;0)\right]\right].$ (89)
Recall that, strictly speaking, for $F(-\mathrm{i}\lambda;0)$ in the second line of (88) to be well-defined, we need $\mathrm{i}\lambda$ to lie within the strip of analyticity, i.e., $\lambda$ must satisfy $\zeta_{+}(0)<\lambda<\zeta_{*}(0)$. Thus (88) is valid for $\lambda\in(-\zeta_{+}(0),\infty)$.
Summarizing, we showed that for $\lambda>-\zeta_{*}(0)$ the function $\hat{S}_{N}(s\,|\,\lambda)$ in the $s$-plane has a simple pole at $s=1$ and a branch point at $s=s_{*}$ with no other singularities for $\left|s\right|<s_{*}$. The schematic illustration is shown in Fig. 9. We emphasize that if $\lambda<-\zeta_{*}(0)$, then additional singularities may arise from the prefactor in (88). However, these singularities have no effect on the large-$n$ behavior of $S_{N}(n\,|\,X_{0})$, as will be explained shortly.
Figure 9. Schematic representation of the analytic structure of the $\hat{S}_{N}(n\,|\,\lambda)$ in the $s$-plane in the survival regime for $\lambda>-\zeta_{*}(0)$. There is a pole at $s=1$ and a branch point at $s_{*}$ defined by (86). In the unshaded region representations (82) and (88) are equivalent. In the shaded region they differ, with (88) providing a proper analytic continuation.
Asymptotic behavior of $S_{N}(n\,|\,X_{0})$
The standard result in the context of analytic combinatorics [62] is that the large-$n$ behavior of the survival probability $S_{N}(n\,|\,X_{0})$ is governed by the singularities of its generating function closest to the origin. To find this generating function, we formally invert the Laplace transform with respect to $\lambda$ and find
$\sum_{n=0}^{\infty}s^{n}S_{N}(n\,|\,X_{0})=\frac{1}{2\pi\mathrm{i}}\int_{\gamma-\mathrm{i}\infty}^{\gamma+\mathrm{i}\infty}\mathrm{d}\lambda\,e^{\lambda X_{0}}\hat{S}_{N}(s\,|\,\lambda),$ (90)
where the integration is performed along the vertical contour that lies to the right of all possible singularities of $\hat{S}_{N}(s\,|\,\lambda)$. For example, we can choose $\gamma$ such that $\gamma>-\zeta_{*}(0)$. With this choice, due to (88), the integrand $S_{N}(s\,|\,\lambda)$ has a pole at $s=1$ and a branch point at $s=s_{*}$. The pole corresponds to the asymptotic value of $S_{N}(n\,|\,X_{0})$ and the branching point at $s_{*}$ determines the rate at which this value is approached.
First and foremost, the residue at $s=1$ determines the asymptotic value of $S_{N}(n\,|\,X_{0})$:
$\lim_{n\to\infty}\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}S_{N}(n\,|\,X_{0})=-\operatorname*{Res}_{s=1}\left[\hat{S}_{N}(s\,|\,\lambda)\right].$ (91)
This limit is nothing but $\hat{S}_{\infty}(\lambda)$, and computing the residue yields
$\hat{S}_{\infty}(\lambda)=\left\{\begin{aligned} &\frac{1}{\lambda}\;\mathcal{G}_{N}(1\,|\,\lambda),\quad\lambda>-\zeta_{*}(0),\\ &\frac{1}{\lambda}\frac{\mathcal{G}_{N}(1\,|\,\lambda)}{1-F\left(-\mathrm{i}\lambda;0\right)},\quad\lambda\in(-\zeta_{+}(0),-\zeta_{*}(0)).\end{aligned}\right$. (92)
Although the representation (92) is slightly different from (69) and (73) obtained in Sec. 4, their equivalence can be established by shifting the contour of integration in $\mathcal{G}_{N}(1\,|\,\lambda)$ as in (89) from $\mathcal{C}_{*}$ back to the real line. Since the detailed analysis of $\hat{S}_{\infty}(\lambda)$ and $S_{\infty}(X_{0})$ was already performed in Sec. 4, here we focus on the corrections for $S_{N}(n\,|\,X_{0})$.
The first subleading correction is given by the singularity in the $s$-plane closest to the origin. In the present case this is a branch cut at $s_{*}$, hence
$\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\left[S_{N}(n\,|\,X_{0})-S_{\infty}(X_{0})\right]\underset{n\to\infty}{\asymp}s_{*}^{-n}.$ (93)
Recall that $s_{*}$ as in (86) does not depend on $\lambda$, hence the $n\to\infty$ behavior of the Laplace transform (93) translates into the $n\to\infty$ behavior of $S_{N}(n\,|\,X_{0})$, namely,
$S_{N}(n\,|\,X_{0})-S_{\infty}(X_{0})\underset{n\to\infty}{\asymp}\exp\left[-\frac{n}{\xi_{n}(\alpha)}\right],$ (94)
with the decay rate given by
$\xi_{n}(\alpha)=\frac{1}{\log s_{*}},\qquad\frac{1}{s_{*}}=\min_{k_{I}}F(\mathrm{i}k_{I};0).$ (95)
5.2. Absorption regime
We now proceed to the absorption regime in which $\mu_{1}(0)<0$ and the process eventually crosses the origin, $S_{\infty}(X_{0})=0$. The analysis is very similar to the one we have just performed in the survival regime. The main difference is that now $\zeta_{*}(0)<0$, and the contour of integration is shifted to the lower part of the $k$-plane. As a result, there is no contribution from the pole at $k=\mathrm{i}\epsilon$, hence, no pole at $s=1$ for $\hat{S}_{N}(s\,|\,\lambda)$, and the analytic continuation reads
$\alpha>\alpha_{c}:\quad\hat{S}_{N}(s\,|\,\lambda)=\left\{\begin{aligned} &\frac{1}{\lambda}\;\mathcal{G}_{N}(s\,|\,\lambda),\quad\lambda>-\zeta_{*}(0),\\ &\frac{1}{\lambda}\frac{\mathcal{G}_{N}(s\,|\,\lambda)}{1-sF\left(-\mathrm{i}\lambda;0\right)},\quad\lambda<-\zeta_{+}(0),\end{aligned}\right$. (96)
where $\mathcal{G}_{N}(s\,|\,\lambda)$ is again given by (89).
Repeating the same arguments as for the survival regime we find that
$S_{N}(n\,|\,X_{0})\underset{n\to\infty}{\asymp}\exp\left[-\frac{n}{\xi_{n}(\alpha)}\right]$ (97)
with the decay rate given by (95).
5.3. Critical point
Finally, we analyze the behavior of $S_{N}(n\,|\,X_{0})$ at the critical point. Specifically we show that as the drift approaches the critical value the <<correlation length>> $\xi_{n}(\alpha)$ diverges resulting in the polynomial decay of the survival probability $S_{N}(n\,|\,X_{0})$ as $n\to\infty$. Additionally, we show that there is a nontrivial limiting behavior as $X_{0}\to\infty$ and $n\to\infty$ with fixed ratio $\frac{X_{0}}{\sqrt{n}}$ and compute the corresponding scaling function.
5.3.1. Divergence of the <<correlation length>>
With the notation (12), the expression (95) for the <<correlation length>> $\xi_{n}(\alpha)$ reads
$\frac{1}{\xi_{n}(\alpha)}=-\log\left\langle e^{-(M-\alpha t)\zeta_{*}}\right\rangle,$ (98)
where $\mathrm{i}\zeta_{*}\equiv\mathrm{i}\zeta_{*}(0)$ is the location of the minimum of $F(k;0)$ along the imaginary axis in the $k$-plane. If $\alpha=\alpha_{c}$, then this minimum is located exactly at the origin, $\zeta_{*}|_{\alpha=\alpha_{c}}=0$, which implies that $\frac{1}{\xi_{n}(\alpha_{c})}=0$; that is $\xi_{n}(\alpha)$ diverges as the strength of the drift approaches its critical value.
To quantify the divergence of $\xi_{n}(\alpha)$ as $\alpha\to\alpha_{c}$ more precisely, we expand (98) in series using the fact that $\zeta_{*}|_{\alpha=\alpha_{c}}=0$. The expansion yields
$\frac{1}{\xi_{n}(\alpha)}\underset{\alpha\to\alpha_{c}}{\sim}-\log\left\langle\vphantom{\frac{}{}}\right.1-(\alpha-\alpha_{c})(M-\alpha_{c}t)\left.\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right|_{\alpha=\alpha_{c}}\\ +\frac{(\alpha-\alpha_{c})^{2}}{2}\left(\left.2t\,\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}+(M-\alpha_{c}t)^{2}\left(\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right)^{2}-(M-\alpha_{c}t)\frac{\mathrm{d}^{2}\zeta_{*}}{\mathrm{d}\alpha^{2}}\right|_{\alpha=\alpha_{c}}\right)\left.\vphantom{\frac{}{}}\right\rangle$. (99)
When taking the average in the logarithm, the term with second derivative of $\zeta_{*}$ disappears and the expansion depends only on the first derivative. Specifically, we obtain
$\frac{1}{\xi_{n}(\alpha)}\underset{\alpha\to\alpha_{c}}{\sim}-\log\left[1+\frac{(\alpha-\alpha_{c})^{2}}{2}\left(\left.2\langle t\rangle\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}+\left\langle(M-\alpha_{c}t)^{2}\right\rangle\left(\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right)^{2}\right|_{\alpha=\alpha_{c}}\right)\right],$ (100)
or, after a simple calculation,
$\frac{1}{\xi_{n}(\alpha)}\underset{\alpha\to\alpha_{c}}{\sim}-\frac{(\alpha-\alpha_{c})^{2}}{2}\left(\left.2\langle t\rangle\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}+\left\langle(M-\alpha_{c}t)^{2}\right\rangle\left(\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right)^{2}\right|_{\alpha=\alpha_{c}}\right).$ (101)
The only thing left is to find the derivative of $\zeta_{*}$ at the critical point. Recall that $\zeta_{*}$ is the position of the minimum, hence it satisfies
$\left\langle(M-\alpha t)e^{-\zeta_{*}(M-\alpha t)}\right\rangle=0.$ (102)
By taking the derivatives with respect to $\alpha$ we can iteratively build the expansion of $\zeta_{*}(\alpha)$. Using that $\zeta_{*}|_{\alpha=\alpha_{c}}=0$ we obtain
$\left\langle(M-\alpha_{c}t)-(\alpha-\alpha_{c})\left(t+(M-\alpha_{c}t)^{2}\left.\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right|_{\alpha=\alpha_{c}}\right)\right\rangle=0,$ (103)
hence
$\left.\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right|_{\alpha=\alpha_{c}}=-\frac{\left\langle t\right\rangle}{\left\langle(M-\alpha_{c}t)^{2}\right\rangle}.$ (104)
Finally, substituting (104) into (101) we get
$\frac{1}{\xi_{n}(\alpha)}\underset{\alpha\to\alpha_{c}}{\sim}\frac{(\alpha-\alpha_{c})^{2}}{2}\frac{\left\langle t\right\rangle^{2}}{\left\langle(M-\alpha_{c}t)^{2}\right\rangle},\qquad\xi_{n}(\alpha)\underset{\alpha\to\alpha_{c}}{\sim}\frac{2}{\left\langle t\right\rangle^{2}}\frac{\left\langle(M-\alpha_{c}t)^{2}\right\rangle}{\left(\alpha-\alpha_{c}\right)^{2}}.$ (105)
This is exactly the behavior stated in (18).
5.3.2. Algebraic decay of $S_{N}(n\,|\,X_{0})$
If $\alpha=\alpha_{c}$, then $s_{*}=1$ and $\xi_{n}(\alpha_{c})=\infty$, hence there is no exponential decay of $S_{N}(n\,|\,X_{0})$ as $n\to\infty$. On the level of the analytic structure of $\hat{S}_{N}(s\,|\,\lambda)$ in the $s$-plane this divergence is manifested by the coincidence of the pole and the branching point, as both of them are located at $s=1$. The large $n$ behavior of $S_{N}(n\,|\,X_{0})$ is governed by the behavior of $\hat{S}_{N}(s\,|\,\lambda)$ in the vicinity of $s=1$.
To construct this expansion we again resort to the representation (62) for $\phi^{\pm}(\lambda;\rho,s)$. Substituting (64) in (78) yields
$\mathrm{Re}(\lambda)>0:\quad\hat{S}_{N}(s\,|\,\lambda)=\frac{1}{s\,\lambda}\frac{2}{\left\langle(M-\alpha_{c}t)^{2}\right\rangle}\frac{1}{k_{-}^{*}(0,s)\left(\lambda+k_{+}^{*}(0,s)\right)}e^{I(\lambda;0,s)},$ (106)
with $k_{\pm}^{*}(\rho,s)$ given by (61). Expanding (106) as $s\to 1$ we obtain
$\mathrm{Re}(\lambda)>0:\quad\hat{S}_{N}(s\,|\,\lambda)\underset{s\to 1}{\sim}\frac{1}{\sqrt{1-s}}\frac{1}{\lambda^{2}}\frac{1}{\sqrt{\frac{1}{2}\langle(M-\alpha_{c}t)^{2}\rangle}}e^{I(\lambda;0,1)},$ (107)
where $I(\lambda;\rho,s)$ is given by (65). The singular behavior (107) suggests that as $n\to\infty$ the survival probability $S_{N}(n\,|\,X_{0})$ decays as $\frac{1}{\sqrt{n}}$. Specifically, we have
$S_{N}(n\,|\,X_{0})\underset{n\to\infty}{\sim}\frac{1}{\sqrt{\pi n}}\,U(X_{0}),$ (108)
with the function $U(X_{0})$ defined through its Laplace transform $\hat{U}(\lambda)$ as
$\mathrm{Re}(\lambda)\geq 0:\quad\hat{U}(\lambda)\equiv\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}U(X_{0})=\frac{1}{\lambda^{2}}\frac{1}{\sqrt{\frac{1}{2}\left\langle(M-\alpha_{c}t)^{2}\right\rangle}}e^{I(\lambda;0,1)}.$ (109)
Now we use (109) to extract the asymptotic behaviors of $U(X_{0})$ for $X_{0}\to 0$ and $X_{0}\to\infty$ by studying the behavior of $\hat{U}(\lambda)$.
The analysis is similar to the one performed in Sec. 4 for the survival probability $S_{\infty}(X_{0})$. There is, however, a major difference. According to (74) the Laplace transform $\hat{S}_{\infty}(\lambda)$ has a singularity of the order $1/\lambda$, which yields a limiting behavior $S_{\infty}(X_{0})\to 1$ as $X_{0}\to\infty$. Extracting the subleading corrections requires a proper analytic continuation to $\mathrm{Re}(\lambda)<0$, which eventually results in (76). At the same time, as is clear from (109), when $\lambda\to 0$ the Laplace transform $\hat{U}(\lambda)$ of $U(X_{0})$ behaves as $1/\lambda^{2}$. Consequently, the leading order behavior for $X_{0}\to\infty$ can be extracted solely from the $\lambda\to 0$ expansion. While the continuation to $\mathrm{Re}(\lambda)<0$ can be straightforwardly performed utilizing the contour deformation in the $k$-plane, to simplify the discussion we do not present it here, especially since the resulting corrections are exponentially small as $X_{0}\to\infty$.
The asymptotic expansions of $U(X_{0})$ as $X_{0}\to 0$ and $X_{0}\to\infty$ are governed by the behaviors of $\hat{U}(\lambda)$ as $\lambda\to\infty$ and $\lambda\to 0$ respectively. The main technical challenge lies in expanding $I(\lambda;0,1)$. We give the detailed derivation in Appendix A, and here present the results. Specifically, substituting expansion (201) into (109) yields, for $X_{0}\to\infty$,
$U(X_{0})\underset{X_{0}\to\infty}{\sim}\frac{1}{\sqrt{\frac{1}{2}\langle(M-\alpha_{c}t)^{2}\rangle}}\left(X_{0}-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\;\frac{1}{k^{2}}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha_{c}t)^{2}\rangle}\right),$ (110)
and similarly due to (207) for $X_{0}\to 0$ we obtain
$U(X_{0})\underset{X_{0}\to 0}{\sim}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha_{c}t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{X_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\log\left|1-F(k;0)\right|\right).$ (111)
We stress that expressions (110) and (111) are explicit for any given distributions $p(t)$ and $q(M)$ that have finite moments and smooth densities.
Before proceeding to the analysis of the survival probability $S_{T}(\tau\,|\,X_{0})$ two remarks are in order. First, the asymptotic expansions (110) and (111) are derived at the critical point $\alpha=\alpha_{c}$, under the assumption that the limit $n\to\infty$ is taken first. In this case, the survival probability $S_{N}(n\,|\,X_{0})$ decays algebraically as $1/\sqrt{n}$, with a prefactor depending on the initial position $X_{0}$. A natural question is, then, what happens when both $n$ and $X_{0}$ tend to infinity simultaneously. This leads to a non-trivial scaling limit, which is essentially equivalent to the Brownian motion.
Second, there exists a special case in which the jump distribution $q(M)$ and the waiting-time distribution $p(t)$ are tuned so that the effective random walk becomes symmetric. Although this situation is restrictive, it provides a valuable benchmark, since the survival probability of symmetric random walks on the half-line has been studied extensively. In this case, our expressions (108), (109), (110) and (111) should recover the results of [51].
Scaling limit
At the critical point $\alpha=\alpha_{c}$, the effective random walk consists of independent identically distributed jumps with zero mean and finite variance. In this situation it is natural to expect convergence to Brownian motion in the limit where both $X_{0}$ and $n$ tend to infinity with a fixed ratio $z=X_{0}/\sqrt{n}$. Accordingly, we anticipate that
$S_{N}(n\,|\,X_{0})\sim V\!\left(z=\frac{X_{0}}{\sqrt{n}}\right),\qquad n\to\infty,\quad X_{0}\to\infty,$ (112)
with a scaling function $V(z)$. Since the scaling limit corresponds to large $n$ and large $X_{0}$, $V(z)$ can be obtained from the expansion of $\hat{S}_{N}(s\,|\,\lambda)$ as $s\to 1$ and $\lambda\to 0$, while respecting the scaling $X_{0}\sim\sqrt{n}$.
To make the connection between the parameters $\lambda$ and $s$ consistent with the scaling relation between $X_{0}$ and $n$, we first change the integration variable in (77) to $z=X_{0}/\sqrt{n}$. This gives
$\hat{S}_{N}(s\,|\,\lambda)=\sum_{n=0}^{\infty}e^{n\log s}\frac{1}{\sqrt{n}}\int_{0}^{\infty}\mathrm{d}z\,e^{-\lambda z\sqrt{n}}S_{N}\left(n\,|\,z\sqrt{n}\right).$ (113)
Next, we introduce the variable y=n(1s)y=n(1-s). In the limit $s\to 1$, the sum over $n$ can be replaced by an integral over $y$, while the survival probability $S_{N}(n\,|\,z\sqrt{n})$ can be replaced by its scaling form $V(z)$,
$\hat{S}_{N}(s\,|\,\lambda)=\frac{1}{(1-s)^{\frac{3}{2}}}\int_{0}^{\infty}\mathrm{d}y\,e^{y\frac{\log s}{1-s}}\frac{1}{\sqrt{y}}\int_{0}^{\infty}\mathrm{d}z\,e^{-\frac{\lambda}{\sqrt{1-s}}z\sqrt{y}}\;V(z).$ (114)
Representation (114) clearly indicates that the correct scaling relation between $\lambda$ and $s$ is $\lambda\sim\sqrt{1-s}$. Introducing the rescaled parameter $u=\lambda/\sqrt{1-s}$, we obtain
$\hat{S}_{N}\left(s\,|\,u\,\sqrt{1-s}\right)\underset{s\to 1}{\sim}\frac{1}{(1-s)^{\frac{3}{2}}}\int_{0}^{\infty}\mathrm{d}y\,e^{-y}\frac{1}{\sqrt{y}}\int_{0}^{\infty}\mathrm{d}z\,e^{-u\,z\sqrt{y}}\;V(z).$ (115)
On the other hand, the same asymptotic behavior can be obtained directly from (106). Recall that $I(\lambda;0,s)$ remains regular as $s\to 1$ and $\lambda\to 0$, so the singularity originates solely from the prefactor. This yields
$\hat{S}_{N}\left(s\,|\,u\,\sqrt{1-s}\right)=\frac{1}{(1-s)^{3/2}}\frac{1}{u\left(u+\sqrt{\frac{1}{2}\left\langle(M-\alpha_{c}t)^{2}\right\rangle}\right)}.$ (116)
Comparing (115) with (116) results in an integral equation for the scaling function $V(z)$:
$\int_{0}^{\infty}\mathrm{d}y\,e^{-y}\frac{1}{\sqrt{y}}\int_{0}^{\infty}\mathrm{d}z\,e^{-u\,z\sqrt{y}}\;V(z)=\frac{1}{u\left(u+\sqrt{\frac{1}{2}\left\langle(M-\alpha_{c}t)^{2}\right\rangle}\right)}.$ (117)
This is precisely the equation studied in detail in [51] in the context of survival probabilities of one-dimensional random walks. Its solution reads
$S_{N}(n\,|\,X_{0})\sim\;\mathrm{erf}\left(\frac{X_{0}}{\sqrt{2n\left\langle(M-\alpha_{c}t)^{2}\right\rangle}}\right),\qquad n\to\infty,\quad\frac{X_{0}}{\sqrt{n}}\text{\penalty 10000\ --- fixed}.$ (118)
It is worth emphasizing that the scaling behavior (118) does not rely on the specific structure of underlying effective random walk beyond the assumptions that it has zero mean, finite moments, and smooth density. In particular, the same functional form of the scaling limit was previously obtained in [51] for any symmetric random walk with finite variance. This essentially reflects the convergence of a random walk to a Brownian motion, which is rather a robust property, insensitive to microscopic details.
Symmetric distribution
Finally, let us consider the case where the effective random walk not only has a zero mean but also is symmetric. Since it requires fine-tuning of the distributions of the jump amplitudes $q(M)$ and inter-jump intervals $p(t)$, such a case is indeed very special. At the same time, the problem of survival probability of symmetric random walks was studied in great detail in [51]. In particular, the behavior of the survival probability of the symmetric random walk with light-tailed increment distribution was shown to be (in our notation) given by
$S_{N}\left(n\,|\,X_{0}\right)\underset{n\to\infty}{\sim}\frac{1}{\sqrt{\pi n}}U_{\text{sym}}(X_{0}),$ (119)
with the function $U_{\text{sym}}(X_{0})$ defined through the Laplace transform as
$\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}U_{\text{sym}}\left(X_{0}\right)=\frac{1}{\lambda}\exp\left[-\frac{\lambda}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\frac{1}{\lambda^{2}+k^{2}}\log\left[1-F(k;0)\right]\right].$ (120)
At first glance the representation (120) may seem to differ from (109), but they are actually the same. Note that the imaginary part of the integrand in (65) is odd, while the real part is even, hence by the parity argument, we can keep only the real part. Since for a symmetric random walk F(k;0)=F(k;0)F(k;0)=F(-k;0) we obtain
$I(\lambda;0,1)=-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\frac{\lambda}{\lambda^{2}+k^{2}}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha t)^{2}\rangle}$ (121)
Then we extract the term with $\log[1-F_{2}(k;0)]$ by using the identity (which can be verified by standard residue calculus)
$\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\frac{\lambda}{\lambda^{2}+k^{2}}\log\left[\frac{1}{2}\mu_{2}(0)k^{2}\right]=\frac{1}{2}\log\left[\frac{1}{2}\mu_{2}(0)\lambda^{2}\right].$ (122)
Substituting then (121) and (122) into (109) we see that it now coincides with (120).
6. Survival probability $S_{T}(\tau\,|\,X_{0})$
Let us now study the properties of $S_{T}(\tau\,|\,X_{0})$, the probability that the process remains positive up to time $\tau$. Specifically, we focus the analytic properties of the Laplace transform of the survival probability
$\hat{S}_{T}(\rho\,|\,\lambda)\equiv\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\int_{0}^{\infty}\mathrm{d}\tau\,e^{-\rho\tau}S_{T}(\tau\,|\,X_{0})$ (123)
in the $\rho$-plane, from which we extract the large $\tau$ behavior of $S_{T}(\tau\,|\,X_{0})$ stated in Sec. 2.1.
First, representing (123) in terms of $\hat{Q}(\rho,s\,|\,\lambda)$ as
$\hat{S}_{T}(\rho\,|\,\lambda)=\frac{1}{\rho}\left.\left(\frac{1}{\lambda}-\hat{Q}\left(\rho,s\,|\,\lambda\right)\right)\right|_{s=1},$ (124)
and using the representation (8) we obtain
$\hat{S}_{T}(\rho\,|\,\lambda)=\lim_{\epsilon\to 0}\frac{1}{\rho}\left[\frac{1}{\lambda}-\frac{1}{\lambda+\frac{\rho}{\alpha}}+\frac{1-c(\rho)}{\lambda+\frac{\rho}{\alpha}}\;\phi^{-}_{\text{a.c}}\left(\lambda+\frac{\rho}{\alpha};\rho,1\right)\phi^{+}_{\text{a.c.}}\left(\epsilon;\rho,1\right)\right].$ (125)
The subsequent derivation follows along the lines of Sec. 5. That is, we look at the integrand (125), move the contour of integration from the real line to $\mathcal{C}_{*}(\rho)$ defined by $\mathrm{i}k=\zeta_{*}(\rho)$ and use the resulting expression for the analytic continuation. The main difference from the continuation of $\hat{S}_{N}(s\,|\,\lambda)$ is that now the new contour $\mathcal{C}_{*}(\rho)$ depends on $\rho$, hence, it moves when $\rho$ changes. This gives rise to more technically involved derivations although the general strategy remains the same. For this reason, below we will be concentrating mostly on the details that differ from $\hat{S}_{N}(s\,|\,\lambda)$.
In addition, in the interest of simplicity, we perform the analytic continuation with respect to $\rho$ only along the real line, eventually finding the pole at $\rho=0$ (determining the survival probability $S_{\infty}(X_{0})$) and a branch point at $\rho=\rho_{*}<0$ (determining the exponential decay of $S_{T}(\tau\,|\,X_{0})$). Since $S_{T}(\tau\,|\,X_{0})$ is essentially a cumulative function of the probability distribution, there are no singularities with $\mathrm{Re}(\rho)>\rho_{*}$ apart from $\rho=0$.
In the following, we separately analyze three cases: the survival regime $\langle M-\alpha t\rangle>0$, the absorption regime $\langle M-\alpha t\rangle<0$, and the critical point $\langle M-\alpha t\rangle=0$.
6.1. Survival regime
When the drift is weak, $\langle M-\alpha t\rangle>0$, the process has a finite probability of never crossing the origin, so that $S_{\infty}(X_{0})>0$. By analyzing the analytic structure of $\hat{S}_{T}(\rho\,|\,\lambda)$ for $\rho<0$ we show that $S_{T}(\tau\,|\,X_{0})$ approaches its asymptotic value exponentially fast with the rate given by (14).
Analytic continuation in the $\rho$-plane
Let us first assume that $\mathrm{Re}\!\left(\lambda+\tfrac{\rho}{\alpha}\right)>0$. In this case according to (55) we have
$\mathrm{Re}\left(\lambda+\frac{\rho}{\alpha}\right)>0:\qquad\phi_{\text{a.c.}}^{\pm}\left(\lambda+\frac{\rho}{\alpha};\rho,1\right)=\phi^{\pm}\left(\lambda+\frac{\rho}{\alpha};\rho,s\right),$ (126)
and hence from (125) we obtain
[…]
The nontrivial analytic structure is entirely contained in the exponential factor. By this stage, the reader has likely developed enough intuition from the previous analysis and may anticipate the next steps. As before, we deform the contour of integration from the real axis to a horizontal contour $\mathcal{C}_{*}(\rho)$ with $\mathrm{Im}(k)=\zeta_{*}(\rho)$, which provides the analytic continuation required in the survival regime.
Now we need to determine whether the deformed contour $\mathcal{C}_{*}(\rho)$ lies above or below the real axis, or, equivalently, to establish the sign of $\zeta_{*}(\rho)$. As argued in Sec. 3.1, this is fixed by the sign of $\mu_{1}(\rho)$. Recall from (29) that
$\mu_{1}(\rho)=\frac{1}{c(\rho)}\int_{0}^{\infty}\mathrm{d}M\,q(M)\,e^{-\rho\frac{M}{\alpha}}\,\bigl(M-\alpha\langle t\rangle\bigr).$ (128)
If $\mu_{1}(0)>0$, then for $\rho<0$ the sign remains unchanged, so that $\zeta_{*}(\rho)>0$ and the contour $\mathcal{C}_{*}(\rho)$, although moving with $\rho$, remains in the upper half-plane.
The analytic structure of the integrand in the $k$-plane is similar to that of (82); the logarithm is analytic within the strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$ and has branch cuts starting at $k=\mathrm{i}\zeta_{1,2}(\rho)$, where $\zeta_{1,2}(\rho)\equiv\zeta_{1,2}(\rho,1)$ are real solutions of (39) with $s=1$ (see Figs. 4 and 5). The term in parentheses provides two simple poles at $k=\mathrm{i}\epsilon$ and $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$. The schematic representation of the analytic structure is given in Fig. 10.
Figure 10. Schematic representation of the analytic structure of the integrand in (127) in the complex $k$-plane within the survival regime. The integrand features two branch points at $k=\mathrm{i}\zeta_{1,2}(\rho)$ and two poles at $k=\mathrm{i}\epsilon$ and $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$. Upon continuation to $\rho<0$, the branch points pinch the shifted contour $\mathcal{C}_{*}(\rho)$ inducing a singularity at $\rho=\rho_{*}$ defined by (130) in the $\rho$-plane.
Therefore, when shifting the contour, the residue at $k=\mathrm{i}\epsilon$ must be accounted for. Specifically, the residue induces a factor $(1-c(\rho)F(\mathrm{i}\epsilon;\rho))^{-1}$. Once the contour is shifted, taking the limit $\epsilon\to 0$ is straightforward, yielding
[…]
We now continue (129) to $\rho<0$. To determine the rightmost singularity, and thereby the asymptotic behavior for large $\tau$, note that (129) can be used as an analytic continuation up to the value $\rho_{*}<0$ at which $\zeta_{1}(\rho)$ and $\zeta_{2}(\rho)$ pinch the contour of integration, i.e.,
$\rho_{*}:\quad\zeta_{1}(\rho_{*};1)=\zeta_{2}(\rho_{*};1)=\zeta_{*}(\rho_{*}).$ (130)
This is essentially the same condition as (85). The coincidence of the branch points implies that at $\rho=\rho_{*}$, equation (39) has only one real solution, which gives an implicit expression for $\rho_{*}$:
$\rho_{*}:\quad\frac{1}{c(\rho_{*})}=\min_{k_{I}}F\left(\mathrm{i}k_{I};\rho_{*}\right).$ (131)
The case $\mathrm{Re}(\lambda+\frac{\rho}{\alpha})<0$ is treated similarly. The difference is that: (i) the analytic continuation in (126) includes a prefactor; (ii) the contour shift from the real line to $\mathcal{C}_{*}(\rho)$ picks up a residue from the pole at $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$. These two contributions cancel, so that the analytic structure in the $\rho$-plane remains unchanged.
Finally, at some point $\rho_{0}$ defined by $\mathrm{Re}(\lambda+\frac{\rho_{0}}{\alpha})>\zeta_{*}(\rho_{0})$, the pole at $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$ traverses the contour $\mathcal{C}_{*}(\rho)$ giving rise to an additional prefactor in front of the integral. As a result, for $\rho_{*}<\rho<\tilde{\rho}$ the analytic continuation of $\hat{S}_{T}(\rho\,|\,\lambda)$ reads
[…]
with
$\mathcal{G}_{T}(\rho\,|\,\lambda)=\exp\left[-\frac{1}{2\pi}\int_{\mathcal{C}_{*}(\rho)}\mathrm{d}k\,\left(\frac{1}{\mathrm{i}k}+\frac{1}{\lambda+\frac{\rho}{\alpha}-\mathrm{i}k}\right)\log\left[1-c(\rho)F(k;\rho)\right]\right].$ (133)
Recall the analogous continuation (88) and (89).
Asymptotic behavior of $S_{T}(\tau\,|\,X_{0})$
Formally inverting the Laplace transform with respect to $\lambda$ we obtain
$\int_{0}^{\infty}\mathrm{d}\tau e^{-\rho\tau}S_{T}(\tau\,|\,X_{0})=\frac{1}{2\pi\mathrm{i}}\int_{\gamma-\mathrm{i}\infty}^{\gamma+\mathrm{i}\infty}\mathrm{d}\lambda\,e^{\lambda X_{0}}\hat{S}_{T}(\rho\,|\,\lambda).$ (134)
Choosing $\gamma$ to be sufficiently large, say $\gamma>-\zeta_{*}(\rho_{*})-\frac{\rho_{*}}{\alpha}$, we see that singularities of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the $\rho$-plane are a pole at $\rho=0$ and the branch point at $\rho=\rho_{*}$.
The pole at $\rho=0$ corresponds to the survival probability $S_{\infty}(X_{0})$. Computing the residue at $\rho=0$ yields
$\mathrm{Re}(\lambda)>0:\quad\hat{S}_{\infty}(\lambda)=\frac{1}{\lambda}\exp\left[-\frac{1}{2\pi}\int_{\mathcal{C}_{*}(0)}\!\mathrm{d}k\left(\frac{1}{\mathrm{i}k}+\frac{1}{\lambda-\mathrm{i}k}\right)\log\left[1-F(k;0)\right]\right].$ (135)
This coincides with (92), which originates from the pole at $s=1$ of $\hat{S}_{N}(s\,|\,\lambda)$. We have already analyzed this expression in detail; here it is treated merely as a consistency check, and we need not analyze it further.
The branch point at $\rho=\rho_{*}$ determines the leading order correction, and implies that the survival probability decays exponentially as
$\alpha<\alpha_{c}:\quad S_{T}(\tau\,|\,X_{0})-S_{\infty}(X_{0})\underset{\tau\to\infty}{\asymp}\exp\left[-\frac{\tau}{\xi_{\tau}(\alpha)}\right],$ (136)
where according to (131), the value of $\rho_{*}$ is determined by solving
$\xi_{\tau}(\alpha)=-\frac{1}{\rho_{*}},\qquad\frac{1}{c(\rho_{*})}=\min_{k_{I}}F\left(\mathrm{i}k_{I};\rho_{*}\right).$ (137)
This is the result stated in (18) for the survival regime.
6.2. Absorption regime
If the drift is strong and $\langle M-\alpha t\rangle<0$, then the process is expected to eventually cross the origin.
On the level of analytic structure in the $\rho$-plane, the difference with respect to the survival regime is that $\mu_{1}(0)<0$, and hence $\mathcal{C}_{*}(0)$ lies in the lower part of the $k$-plane. Consequently, when shifting the contour, there is no contribution from the pole at $k=\mathrm{i}\epsilon$, and thus
[…]
Notably, the expression above has no pole at $\rho=0$, since $c(0)=1$.
To determine the asymptotic behavior of the survival probability, we analyze the pinching of the contour $\mathcal{C}_{*}(\rho)$ by branch points at $k=\zeta_{1,2}(\rho)$ in the $k$-plane. From (128) it follows that at some $\rho_{c}<0$, the function $\mu_{1}(\rho)$ changes sign, causing the contour $\mathcal{C}_{*}(\rho)$ to shift from the lower to the upper half of the $k$-plane. This could, in principle, alter the asymptotic behavior by changing the dominant singularities, but, in fact, it does not.
If $\rho=\rho_{c}$, then $\mu_{1}(\rho_{c})=0$, and hence the function $F(\mathrm{i}k;\rho_{c})$ reaches its minimum at $k=0$, with $F(0;\rho_{c})=1$. Since $\rho_{c}$ is negative, we have $c(\rho_{c})>1$, and thus the equation (39) has no real solutions. This confirms that when continuing $\rho$ to the negative half-line, the contour pinching occurs before any potential complications caused by the movement of the shifted contour $\mathcal{C}_{*}(\rho)$. In other words, $\rho_{c}<\rho_{*}$, ensuring that the decay rate of the survival probability is indeed determined by $\rho_{*}$.
Repeating now the same arguments as in the survival regime we find that
$\alpha>\alpha_{c}:\qquad S_{T}(\tau\,|\,X_{0})\underset{\tau\to\infty}{\asymp}\exp\left[-\frac{\tau}{\xi_{\tau}(\alpha)}\right],$ (139)
with the rate given by (137). This is the result stated in (18) for the absorption regime.
6.3. Critical point
We conclude the analysis of the survival probability $S_{T}(\tau\,|\,X_{0})$ by examining the critical point where $\langle M-\alpha t\rangle=0$. The analysis follows closely the approach presented in Sec. 5.3 for $S_{N}(n\,|\,X_{0})$. Therefore, we outline only the main steps and present the key results.
6.3.1. Divergence of the <<correlation length>>
We first demonstrate that as the drift strength $\alpha$ approaches the critical value $\alpha_{c}$, the correlation length $\xi_{\tau}(\alpha)$ defined in (137) diverges.
Using the notation (12), the equation (137) for $\rho_{*}$ can be written as
$\left\langle\exp\left[-\rho_{*}\frac{M}{\alpha}-\zeta_{*}(\rho_{*})(M-\alpha t)\right]\right\rangle=1.$ (140)
At the critical point, we have $\mu_{1}(0)=0$, which implies
$\rho_{*}\big|_{\alpha=\alpha_{c}}=0,\qquad\zeta_{*}\big|_{\alpha=\alpha_{c}}=0.$ (141)
Treating both $\rho_{*}$ and $\zeta_{*}$ as functions of $\alpha$ and expanding (140) to first order as $\alpha\to\alpha_{c}$, we obtain
$1=\left\langle 1-(\alpha-\alpha_{c})\left.\left[(M-\alpha_{c}t)\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}-\frac{M}{\alpha_{c}}\frac{\mathrm{d}\rho_{*}}{\mathrm{d}\alpha}\right]\right|_{\alpha=\alpha_{c}}\right\rangle+O\left((\alpha-\alpha_{c})^{2}\right).$ (142)
Taking the average and using $\langle M-\alpha_{c}t\rangle=0$, we find
$\left.\frac{\mathrm{d}\rho_{*}}{\mathrm{d}\alpha}\right|_{\alpha=\alpha_{c}}=0.$ (143)
Further expanding (140) to the second order and examining the coefficient of $(\alpha-\alpha_{c})^{2}$, we obtain the consistency condition
$\left\langle\left.\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\left(2t+(M-\alpha_{c}t)^{2}\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right)-\left(M-\alpha_{c}t\right)\frac{\mathrm{d}^{2}\zeta_{*}}{\mathrm{d}\alpha^{2}}-\frac{\mathrm{d}^{2}\rho_{*}}{\mathrm{d}\alpha^{2}}\frac{M}{\alpha_{c}}\right|_{\alpha=\alpha_{c}}\right\rangle=0,$ (144)
and hence
$\left.\frac{\mathrm{d}^{2}\rho_{*}}{\mathrm{d}\alpha^{2}}\right|_{\alpha=\alpha_{c}}=\frac{\alpha_{c}}{\langle M\rangle}\left.\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\left(2\langle t\rangle+\left\langle(M-\alpha_{c}t)^{2}\right\rangle\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right)\right|_{\alpha=\alpha_{c}}.$ (145)
To complete the calculation, we need the derivative of $\zeta_{*}$ at the critical point. Since $\zeta_{*}$ is the location of the minimum of $F(\mathrm{i}k_{I};\rho_{*})$, it satisfies
$\left\langle(M-\alpha t)\exp\left[-\rho_{*}\frac{M}{\alpha}-\zeta_{*}(M-\alpha t)\right]\right\rangle=0.$ (146)
Expanding as $\alpha\to\alpha_{c}$ gives
$\left.\frac{\mathrm{d}\zeta_{*}}{\mathrm{d}\alpha}\right|_{\alpha=\alpha_{c}}=-\frac{\left\langle t\right\rangle}{\left\langle(M-\alpha_{c}t)^{2}\right\rangle}.$ (147)
Note that due to (141) and (143), in the first order, (146) is equivalent to (102), and hence (147) coincides with (104). Substituting now (147) into (145) yields
$\left.\frac{\mathrm{d}^{2}\rho_{*}}{\mathrm{d}\alpha^{2}}\right|_{\alpha=\alpha_{c}}=-\frac{\alpha_{c}}{\left\langle M\right\rangle}\frac{\langle t\rangle^{2}}{\left\langle(M-\alpha_{c}t)^{2}\right\rangle}$ (148)
which according to (137) translates into the divergence
$\xi_{\tau}(\alpha)\underset{\alpha\to\alpha_{c}}{\sim}\frac{\langle M\rangle}{\alpha_{c}}\frac{2}{\langle t\rangle^{2}}\frac{\langle(M-\alpha_{c}t)^{2}\rangle}{(\alpha-\alpha_{c})^{2}}.$ (149)
Recalling that $\alpha_{c}=\frac{\langle M\rangle}{\langle t\rangle}$ we obtain the result stated in (18).
6.3.2. Algebraic decay of $S_{T}(\tau\,|\,X_{0})$
From (149) it is clear that $\xi_{\tau}(\alpha_{c})=\infty$ and hence there is no exponential decay of the survival probability $S_{T}(\tau\,|\,X_{0})$, and algebraic behavior is expected instead. On the level of analytic structure of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the $\rho$-plane, it corresponds to the coincidence of the branch point and the pole (recall that at the critical point $\rho_{*}=0$). The large $\tau$ behavior of $S_{T}(\tau\,|\,X_{0})$ is therefore governed by the expansion of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the vicinity of $\rho=0$.
Substituting the representation (64) into (124) and expanding the resulting expression for $\rho\to 0$ we obtain
$\mathrm{Re}(\lambda)>0:\quad\hat{S}_{T}(\rho\,|\,\lambda)\underset{\rho\to 0}{\sim}\frac{1}{\sqrt{\rho}}\frac{1}{\lambda^{2}}\sqrt{\frac{\langle t\rangle}{\frac{1}{2}\left\langle(M-\alpha_{c}t)^{2}\right\rangle}}e^{I(\lambda;0,1)},$ (150)
where we used
$c(\rho)=1-\frac{\left\langle M\right\rangle}{\alpha}\rho+O(\rho^{2})=1-\langle t\rangle\rho+O(\rho^{2}).$ (151)
Note that (150) is very similar to (107), and the square root divergence implies that the large $\tau$ behavior is
$S_{T}(\tau\,|\,X_{0})\underset{\tau\to\infty}{\sim}\sqrt{\frac{\langle t\rangle}{\pi\tau}}U(X_{0}),$ (152)
where $U(X_{0})$ is the same function as for the survival probability $S_{N}(n\,|\,X_{0})$. Recall that it is defined through the Laplace transform as in (109), and the asymptotic behaviors for $X_{0}\to\infty$ and $X_{0}\to 0$ are given by (110) and (111) respectively.
As argued in Sec. 5.3, at the critical point, after a proper rescaling the process converges to a Brownian motion, and the survival probability $S_{N}(n\,|\,X_{0})$ admits a scaling form defined by (118). Repeating essentially the same derivation we obtain
$S_{T}(\tau\,|\,X_{0})\sim\;\mathrm{erf}\left(\frac{X_{0}}{\sqrt{2\frac{\tau}{\langle t\rangle}\langle(M-\alpha_{c}t)^{2}\rangle}}\right),\qquad\tau\to\infty,\quad\frac{X_{0}}{\sqrt{\tau}}\text{\penalty 10000\ --- fixed}.$ (153)
With this we conclude the analysis of the survival probabilities.
7. Moments of $\tau$ and $n$
Having established and thoroughly quantified the large $\tau$ and large $n$ behavior of the survival probabilities using the representation (8), the next natural step is to consider other quantities of interest. One important question concerns the moments of the first-passage time $\tau$ and the number of jumps before the first passage $n$ in the absorption regime. In this section we address this question and compute the first two leading terms in the $X_{0}\to\infty$ and $X_{0}\to 0$ expansions for the means and the variances of $\tau$ and $n$.
The Laplace transform of the moments can be obtained by differentiating the function $\hat{Q}(\rho,s\,|\,\lambda)$. For example, the mean values are
$\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\mathbb{E}\left[\tau\,|\,X_{0}\right]=-\left.\frac{\partial}{\partial\rho}\hat{Q}(\rho,s\,|\,\lambda)\right|_{\rho=0,s=1},$ (154)
$\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\mathbb{E}\left[n\,|\,X_{0}\right]=\left.\frac{\partial}{\partial s}\hat{Q}(\rho,s\,|\,\lambda)\right|_{\rho=0,s=1},$ (155)
or, using the representation (64),
$\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\mathbb{E}\left[n\,|\,X_{0}\right]=-\frac{1}{\lambda^{2}}\frac{1}{\langle M-\alpha t\rangle}e^{I(\lambda;0,1)},$ (156)
$\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\mathbb{E}\left[\tau\,|\,X_{0}\right]=\frac{1}{\alpha\lambda^{2}}-\frac{1}{\alpha\lambda^{2}}\frac{\langle M\rangle}{\langle M-\alpha t\rangle}e^{I(\lambda;0,1)},$ (157)
where $I(\lambda;\rho,s)$ is given by
$I(\lambda;\rho,s)=-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\,\left(\frac{1}{\mathrm{i}k}+\frac{1}{\lambda+\frac{\rho}{\alpha}-\mathrm{i}k}\right)\log\frac{1-s\,c(\rho)\,F(k;\rho)}{1-s\,c(\rho)\,F_{2}(k;\rho)}.$ (158)
The representations (156) and (157) however still involve rather complicated integrals, and their practical utility might not be immediately apparent. Especially since we are dealing with a very general class of distributions, and it is unrealistic to expect that the Laplace transforms can be inverted analytically for arbitrary $p(t)$ and $q(M)$.
What can be done instead is to extract the asymptotic behavior of the mean values as $X_{0}\to\infty$ and $X_{0}\to 0$ from the expansions of the Laplace transforms as $\lambda\to 0$ and $\lambda\to\infty$. Typically such expansions can be constructed by expanding the integrand in $I(\lambda;0,1)$ term by term, and the rest becomes a technical exercise. In our case, however, this naive approach results in divergent integrals, necessitating a more careful treatment.
The resolution lies in recognizing that $I(\lambda;\rho,s)$ can be expressed as Mellin convolution integrals, for which powerful asymptotic techniques exist. In this section, we illustrate how such asymptotic expansions can be constructed via Mellin transform techniques, specifically applying the method presented in [63]. We compute the first two terms in the expansion of both the mean and the variance for $\tau$ and $n$ as $X_{0}\to\infty$ and $X_{0}\to 0$ (16 terms in total). This demonstrates that the mapping to an effective random walk, together with the representation (8), provides a powerful tool.
Before proceeding to the technical details, let us make a useful observation. From (156) and (157), the means of $\tau$ and $n$ are related via
$\mathbb{E}\left[\tau\,|\,X_{0}\right]=\frac{X_{0}}{\alpha}+\frac{\langle M\rangle}{\alpha}\mathbb{E}\left[n\,|\,X_{0}\right].$ (159)
Relation (159) can readily be used to recover the asymptotic expansions for the mean value of $\tau$ from those of the mean value of $n$. In other words, instead of 16 terms we need to compute only 12.
Asymptotic expansions of $I(\lambda;\rho,s)$
Computing the $X_{0}\to\infty$ and $X_{0}\to 0$ asymptotic behaviors of the first two moments involves asymptotic expansions of $I(\lambda;\rho,s)$ (for means) and its first derivatives with respect to $\rho$ and $s$ (for second moments of $\tau$ and $n$) as $\lambda\to 0$ and $\lambda\to\infty$. As previously mentioned, simply expanding the integrand in series with respect to $\lambda$ results in divergent integrals. Leaving the verification of this fact to the reader, we present an alternative approach based on Mellin transform techniques.
We first simplify the integral representation. Note that the imaginary part of the integrand in (158) is odd, whereas the real part is even. By parity, we can write
$I(\lambda;\rho,s)=-\frac{1}{\pi}\int_{-\infty}^{\infty}\mathrm{d}k\,\mathrm{Re}\left\{\left(\frac{1}{\mathrm{i}k}+\frac{1}{\lambda+\frac{\rho}{\alpha}-\mathrm{i}k}\right)\log\frac{1-s\,c(\rho)\,F(k;\rho)}{1-s\,c(\rho)\,F_{2}(k;\rho)}\right\}.$ (160)
Let us now focus on $I(\lambda;0,1)$. Setting $\rho=0$ and $s=1$ in (160), rescaling the integration variable $k\mapsto\lambda k$, and separating the real and imaginary parts explicitly, we obtain
$I(\lambda;0,1)=-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{2}+1}\log\left|\frac{1-F(\lambda k;0)}{1-F_{2}(\lambda k;0)}\right|\\ -\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k(k^{2}+1)}\arg\frac{1-F(\lambda k;0)}{1-F_{2}(\lambda k;0)}.$ (161)
Each integral is now of the form
$I(\lambda)=\int_{0}^{\infty}\mathrm{d}k\,f(k)\,h(\lambda k),$ (162)
which are Mellin convolution integrals. There exists a well-developed body of theory regarding the asymptotic analysis of such integrals. Here we employ the method from [63] for its simplicity and directness.
The key result is as follows. Suppose we aim to find $\lambda\to 0$ behavior of (162), where the functions $f(k)$ and $h(k)$ have asymptotic expansions
$f(k)\underset{k\to\infty}{=}\sum_{\ell=0}^{n-1}\frac{a_{\ell}}{k^{\alpha_{\ell}}}+O\left(k^{-\alpha_{n}}\right),\qquad h(k)\underset{k\to 0}{=}\sum_{\ell=0}^{n-1}b_{\ell}k^{\beta_{\ell}}+O(k^{\beta_{n}})$ (163)
with $\alpha_{\ell}-\beta_{j}\neq 1$ for all pairs $(\ell,j)$. Then the asymptotic behavior of (162) reads
$I(\lambda)=\sum_{\ell=0}^{n-1}a_{\ell}\lambda^{\alpha_{\ell}-1}\mathcal{M}[h;1-\alpha_{\ell}]+\sum_{j=0}^{m-1}b_{j}\lambda^{\beta_{j}}\mathcal{M}[f;\beta_{j}+1]+O\left(\lambda^{\beta_{m}}+\lambda^{\alpha_{n}-1}\right),$ (164)
where $\mathcal{M}[g;z]$ is the Mellin transform defined as the integral
$\mathcal{M}[g;z]=\int_{0}^{\infty}k^{z-1}g(k)\mathrm{d}k,$ (165)
or its analytic continuation in the region of $z$ where the integral does not exist.
If there exist pairs $(\ell,j)$ such that $\alpha_{\ell}-\beta_{j}=1$, the asymptotic behavior in (164) must be modified. Although such situations do not arise in our context, we present the necessary modification for completeness. Specifically, for each such pair $(\ell,j)$, the sum of terms
$a_{\ell}\lambda^{\alpha_{\ell}-1}\mathcal{M}[h;1-\alpha_{\ell}]+b_{j}\lambda^{\beta_{j}}\mathcal{M}[f;\beta_{j}+1]$ (166)
is replaced by
$\lambda^{\beta_{j}}\lim_{z\to 0}\left\{\vphantom{\frac{}{}}a_{\ell}\mathcal{M}[h;z+1-\alpha_{\ell}]+b_{j}\mathcal{M}[f;z+1+\beta_{j}]\right\}-a_{\ell}b_{j}\;\lambda^{\beta_{j}}\log\lambda.$ (167)
Furthermore, if $\alpha_{n}-\beta_{m}=1$, then $O\left(\lambda^{\beta_{m}}+\lambda^{\alpha_{n}-1}\right)$ must be changed to $O\left(\lambda^{\beta_{m}}\log\lambda\right)$.
The beauty of (164) together with (167) is that it systematically handles the divergences that would appear if naively expanding (158), converting them into well-defined Mellin transforms.
The final step is to apply (164) to $I(\lambda;0,1)$ and to its derivatives $\partial_{s}I(\lambda;0,1)$ and $\partial_{\rho}I(\lambda;0,1)$. The $\lambda\to 0$ expansions yield the $X_{0}\to\infty$ behavior, while $\lambda\to\infty$ expansions (obtained by reversing the roles of $f$ and $h$) give the $X_{0}\to 0$ behavior. Although the calculations involve considerable algebraic effort, they reduce to systematic application of Taylor expansions and Mellin transform evaluations.
Asymptotic expansions for the mean and the variance
Having established a systematic method to compute the asymptotic expansions of $I(\lambda;0,1)$ and its derivatives, the remaining task is computational. Rather than burden the reader with extensive algebraic details, we defer the computational part to Appendix A and proceed directly to presenting the final results in the form of asymptotic expansions
$\mathbb{E}[n\,|\,X_{0}]\underset{X_{0}\to\infty}{\sim}X_{0}A_{1}+A_{0},\qquad\mathbb{E}[n\,|\,X_{0}]\underset{X_{0}\to 0}{\sim}a_{0}+a_{1}X_{0},$ (168)
$\mathrm{Var}[n\,|\,X_{0}]\underset{X_{0}\to\infty}{\sim}X_{0}B_{1}+B_{0},\qquad\mathrm{Var}[n\,|\,X_{0}]\underset{X_{0}\to 0}{\sim}b_{0}+b_{1}X_{0},$ (169)
$\mathrm{Var}[\tau\,|\,X_{0}]\underset{X_{0}\to\infty}{\sim}X_{0}C_{1}+C_{0},\qquad\mathrm{Var}[\tau\,|\,X_{0}]\underset{X_{0}\to 0}{\sim}c_{0}+c_{1}X_{0}.$ (170)
Note that the asymptotics of $\mathbb{E}[\tau\,|\,X_{0}]$ are recovered from (168) using the relation (159).
The coefficients in (168), (169), and (170) are expressed in terms of the following auxiliary integrals:
$J_{1}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right],$ (171)
$J_{2}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\,\mathrm{Re}\left[\frac{1}{k^{2}}\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right],$ (172)
$J_{3}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Im}\left[\frac{1}{k^{3}}\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right],$ (173)
$K_{1}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\frac{1}{k}\,\mathrm{Im}\left[\frac{1}{1-F(k;0)}-\frac{1}{1-F_{2}(k;0)}\right],$ (174)
$K_{2}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\frac{1}{k^{2}}\,\mathrm{Re}\left[\frac{1}{1-F(k;0)}-\frac{1}{1-F_{2}(k;0)}\right],$ (175)
$D_{1}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\left.\frac{1}{k}\,\mathrm{Im}\left[\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;0)}-\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;0)}\right]\right|_{\rho=0},$ (176)
$D_{2}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\left.\frac{1}{k^{2}}\,\mathrm{Re}\left[\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;0)}-\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;0)}\right]\right|_{\rho=0},$ (177)
$L_{1}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Re}\left[\log(1-F(k;0))\right],$ (178)
$L_{2}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Re}\left[\frac{F(k;0)}{1-F(k;0)}\right],$ (179)
$L_{3}=\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Re}\left[\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;0)}\right].$ (180)
While these integrals appear complex, their key virtue is that they provide explicit expressions valid for arbitrary distributions $p(t)$ and $q(M)$. Given specific distributions, these integrals can be evaluated numerically with standard techniques.
The advantage of the effective random walk approach becomes particularly evident when compared to alternative methods. Traditional renewal equation techniques, while elegant for special cases, face significant challenges when dealing with general distributions and higher-order moments. In contrast, the triple Laplace transform representation (8) provides a systematic computational framework with no conceptual barriers. The method extends naturally to higher moments: computing asymptotic expansions for third or fourth moments would require evaluating second or third derivatives of $I(\lambda;\rho,s)$, leading to additional auxiliary integrals similar to those above. The computational complexity would indeed grow, but the calculation remains tractable through systematic application of Taylor expansions and Mellin transform techniques.
The explicit coefficient expressions presented below serve primarily as a concrete demonstration of this computational power. They illustrate that the effective random walk mapping, combined with our integral representation, transforms what would otherwise be barely tractable asymptotic problems into algebraically intensive, yet very systematic, calculations.
We now give the explicit coefficient expressions. For the mean number of jumps (168), the $X_{0}\to\infty$ coefficients are
$A_{1}=-\frac{1}{\langle M-\alpha t\rangle},\qquad A_{0}=\frac{J_{2}}{\langle M-\alpha t\rangle},$ (181)
while the $X_{0}\to 0$ coefficients are
$a_{0}=\frac{\sqrt{\frac{1}{2}\langle(M-\alpha t)^{2}\rangle}}{\langle M-\alpha t\rangle}e^{-J_{1}},\qquad a_{1}=a_{0}\;L_{1}.$ (182)
For the variance of $n$ the coefficients in the expansion (169) for $X_{0}\to\infty$ are
$B_{1}$ $=-\frac{\langle M^{2}\rangle-\langle M\rangle^{2}+\langle(\alpha t)^{2}\rangle-\langle\alpha t\rangle^{2}}{\langle M-\alpha t\rangle^{3}},$ (183)
$B_{0}$ $=-B_{1}J_{2}-\frac{\langle(M-\alpha t)^{3}\rangle}{3\langle M-\alpha t\rangle^{3}}-\frac{2J_{3}}{\langle M-\alpha t\rangle^{2}}-\frac{2K_{2}}{\langle M-\alpha t\rangle},$
and, similarly for $X_{0}\to 0$ we have:
$b_{0}=a_{0}\left[2K_{1}-a_{0}+\frac{\langle(M-\alpha t)^{2}\rangle}{\langle M-\alpha t\rangle^{2}}\right],$ (184)
$b_{1}=L_{1}\left(a_{0}^{2}-b_{0}\right)-\frac{a_{0}}{\langle M-\alpha t\rangle}+2a_{0}L_{2}.$
Finally, for the variance of $\tau$ given in (170), the $X_{0}\to\infty$ coefficients are
$C_{1}$ $=-\frac{\langle M\rangle^{2}\left[\langle t^{2}\rangle-\langle t\rangle^{2}\right]-\langle t\rangle^{2}\left[\langle M^{2}\rangle-\langle M\rangle^{2}\right]}{\left\langle M-\alpha t\right\rangle^{3}},$ (185)
$C_{0}$ $=-C_{1}J_{2}-\frac{\langle M\rangle^{2}}{3\alpha^{2}}\frac{\langle(M-\alpha t)^{3}\rangle}{\langle M-\alpha t\rangle^{3}}-\frac{\langle M\rangle^{2}}{\alpha^{2}}\frac{2J_{3}}{\langle M-\alpha t\rangle^{2}}+\frac{\langle M\rangle}{\alpha}\frac{2D_{2}}{\langle M-\alpha t\rangle},$
and the $X_{0}\to 0$ coefficients are
$c_{0}=$ $-\frac{\langle M\rangle^{2}}{\alpha^{2}}a_{0}^{2}-2\frac{\langle M\rangle}{\alpha}a_{0}D_{1}+a_{0}\frac{\langle M\rangle\langle M(M-\alpha t)^{2}\rangle}{\alpha^{2}\langle(M-\alpha t)^{2}\rangle}$ (186)
$\qquad+a_{0}\frac{\langle M^{2}\rangle\left[\langle t^{2}\rangle-\langle t\rangle^{2}\right]+\langle t^{2}\rangle\left[\langle M^{2}\rangle-\langle M\rangle^{2}\right]}{\langle M-\alpha t\rangle^{2}},$
$c_{1}=$ $L_{1}\left(a_{0}^{2}\frac{\langle M\rangle^{2}}{\alpha^{2}}-c_{0}\right)-\frac{\langle M\rangle^{2}}{\alpha^{2}}\frac{a_{0}}{\langle M-\alpha t\rangle}-2a_{0}\,L_{3}\frac{\langle M\rangle}{\alpha}.$
Numerical verification
Having derived rather cumbersome expressions for the asymptotic coefficients, it is imperative to test their validness. The first check is to compare them against the exactly solvable case where both $p(t)$ and $q(M)$ are exponentially distributed, for which exact analytical expressions were obtained in [36]. A motivated reader can verify that the asymptotic expressions do coincide with these exact results.
Figure 11. Comparison between the numerical simulations and the asymptotic expansions of the mean values of $\tau$ and $n$ for the distributions (187). The parameters of the model are $\alpha=2$, $\gamma=1$, $\ell=0{.}9$. Black dashed lines correspond to asymptotic expansions for $X_{0}\to\infty$. Blue dash-dotted lines correspond to asymptotic expansions for $X_{0}\to 0$. Red solid lines correspond to the numerical simulation, which was obtained be generating $10^{8}$ trajectories for 5050 values of $X_{0}$ in $[0;2]$
However, testing the asymptotics only against this special case would not be prudent. In the double exponential case, the mean and variance are actually linear functions of $X_{0}$, which means that the asymptotic behaviors for $X_{0}\to\infty$ and $X_{0}\to 0$ are essentially identical. More importantly, this might mislead the reader into believing that such linear behavior is generic. This suggests that we need to test the asymptotic expansions against a more general case of the distributions $p(t)$ and $q(M)$, making numerical simulations the natural choice. Following this idea, we performed Monte Carlo simulations for a specific process where the time intervals are drawn from a half-Gaussian distribution and the jumps follow a uniform distribution:
$q(M)=\sqrt{\frac{2}{\pi}}\gamma\;e^{-\frac{1}{2}\gamma^{2}M^{2}},\qquad p(t)=\begin{cases}\frac{1}{\ell},&0<t\leq\ell,\\ 0,&t>\ell.\end{cases}$ (187)
For this choice we have
$\left\langle M-\alpha t\right\rangle=\frac{1}{\gamma}\sqrt{\frac{2}{\pi}}-\frac{\alpha\ell}{2}.$ (188)
The specific choice (187) illustrates two key features: the support of $p(t)$ is bounded rather than infinite, and this combination appears to have no analytical solution — neither the auxiliary integrals nor the corresponding renewal equations have known closed forms.
By generating trajectories for different initial positions, we computed the sample means and variances of $\tau$ and $n$ and compared them with our theoretical predictions. The comparison in Fig. 11 demonstrates perfect agreement and confirms the predicted asymptotic behavior.
8. Conclusion
In this paper, we applied the mapping of the jump process with drift to an effective random walk for the case where both inter-arrival times and jump amplitudes follow arbitrary light-tailed distributions with smooth densities. Through detailed analysis of the Laplace transform of the joint probability distribution of $\tau$ and $n$, we obtained exact asymptotic results for general distributions.
Our main result is confirming that the intuitive picture obtained in [36] for exactly solvable cases extends to the general setting. We identified three regimes: the survival regime (weak drift) where the process has finite probability of remaining positive indefinitely, the absorption regime (strong drift) where first-passage occurs with certainty, and the critical point separating these regimes. In both survival and absorption regimes, survival probabilities approach their asymptotic values exponentially fast, with explicitly computable decay rates. At the critical point, exponential decay is replaced by algebraic decay with universal scaling behavior described by error functions.
Beyond the decay rates, we derived asymptotic expansions for the mean and variance of both $\tau$ and $n$ in the limits $X_{0}\to\infty$ and $X_{0}\to 0$. All results are expressed in terms of the Fourier transform of the effective random walk, making them applicable to arbitrary distributions. This demonstrates the power of the random walk mapping approach for analyzing first-passage properties beyond exactly solvable cases.
A natural question concerns the case where our assumptions are violated. As mentioned earlier, the smooth density requirement is largely technical and can likely be relaxed without substantial modifications to the analysis. The case of heavy-tailed distributions, where the second moment of the effective random walk is infinite, on the other hand, presents a challenge. Allowing the distributions to be heavy-tailed would alter the behavior of $F(k;\rho)$ in ways that could invalidate our asymptotic analysis. The heavy-tailed case therefore merits detailed investigation. Another important assumption is that $t_{i}$ and $M_{j}$ are independent. It would be interesting to study whether the proposed formalism can be generalized to some extent to situations where correlations are present, as was done for the Sparre-Andersen theorem in [64].
[…]
Acknowledgements
The author is indebted to S. N. Majumdar for his guidance and valuable discussions throughout this work.
Appendix A Asymptotic expansions for the means and variances
This appendix is devoted to the computation of the asymptotic expansions of the means and variances of $n$ and $\tau$. First we extract the asymptotic behavior of $I(\lambda;\rho,s)$ and its two first derivatives using the approach of [63]. A key ingredient is the expression (164) that allows us to construct the asymptotic expansion for the Mellin convolution integrals (162) from the asymptotic behavior of the convoluted functions (163). These results are then used to construct the expansions of the mean and variances given by (168), (169), and (170).
A.1. Expansion of $I(\lambda;0,1)$
We begin with the expansions of $I(\lambda;0,1)$, which appears in the expressions for both the means and variances of $\tau$ and $n$.
Expansion $\lambda\to 0$.
We first construct the $\lambda\to 0$ expansion of $I(\lambda;0,1)$. Recall that according to (161) this integral is given by
$I(\lambda;0,1)=-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,f_{1}(k)h_{1}(\lambda k)-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,f_{2}(k)h_{2}(\lambda k).$ (189)
We recognize that (189) is decomposed into two separate Mellin convolution integrals of the form (162), with the functions $f_{1,2}(k)$ are given by
$f_{1}(k)=\frac{1}{k^{2}+1},\qquad f_{2}(k)=\frac{1}{k(k^{2}+1)},$ (190)
and for the functions $h_{1,2}(k)$ we have
$h_{1}(k)=\,\mathrm{Re}\left[\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right],\quad h_{2}(k)=\,\mathrm{Im}\left[\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right].$ (191)
To apply (164) to the first term we need the asymptotic expansions of $f_{1}(k)$ as $k\to\infty$ and of $h_{1}(k)$ as $k\to 0$. The former is straightforward and the latter follows from expansion (28) of the Fourier transform $F(k;\rho)$ for $k\to 0$, namely
$f_{1}(k)\underset{k\to\infty}{=}\frac{1}{k^{2}}+O\left(\frac{1}{k^{4}}\right),\qquad h_{1}(k)\underset{k\to 0}{=}-\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)}\,k^{2}+O\left(k^{3}\right).$ (192)
Hence in the notation (163) we have
$a_{0}=1,\quad\alpha_{0}=2,\quad\alpha_{1}=4,\quad b_{0}=-\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)},\quad\beta_{0}=2,\quad\beta_{1}=3.$ (193)
Then we apply the general formula (164). Substituting (193) into this general result yields
$\int_{0}^{\infty}\mathrm{d}k\,f_{1}(k)h_{1}(\lambda k)\underset{\lambda\to 0}{=}\lambda\;\mathcal{M}[h_{1};-1]-\lambda^{2}\,\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)}\mathcal{M}[f_{1};-1]+O(\lambda^{3}).$ (194)
To proceed further, we evaluate the Mellin transform of $f_{1}(k)$, which can be computed using standard complex analysis techniques,
$\mathcal{M}\left[f_{1};z\right]=\int_{0}^{\infty}\mathrm{d}k\;k^{z-1}\frac{1}{k^{2}+1}=\frac{\pi}{2}\frac{1}{\sin\frac{\pi z}{2}},\qquad\mathcal{M}[f_{1};-1]=-\frac{\pi}{2}.$ (195)
Since $h_{1}$ explicitly depends on $F(k;0)$ and hence is determined by the exact form of the distributions $p(t)$ and $q(M)$ we keep its Mellin transform in the integral form. The contribution from the first term in (189) then becomes
$\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,f_{1}(k)h_{1}(\lambda k)\underset{\lambda\to 0}{\sim}\frac{\lambda}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{2}}\log\left|\frac{1-F(k;0)}{1-F_{2}(k;0)}\right|+\frac{\lambda^{2}}{12}\frac{\mu_{3}(0)}{\mu_{1}(0)}.$ (196)
We now turn to the second term in (189), for which the convoluted functions are $f_{2}(k)$ and $h_{2}(k)$. The asymptotic analysis proceeds similarly, though the function $h_{2}$ has a higher-order vanishing at $k=0$,
$f_{2}(k)\underset{k\to\infty}{=}\frac{1}{k^{3}}+O\left(\frac{1}{k^{5}}\right),\qquad h_{2}(k)\underset{k\to 0}{=}O\left(k^{3}\right).$ (197)
In the notation (163) we have
$a_{0}=1,\quad\alpha_{0}=3,\quad\alpha_{1}=5,\quad\beta_{0}=3,$ (198)
and hence the asymptotic behavior of the second term reads
$\int_{0}^{\infty}\mathrm{d}k\,f_{2}(k)h_{2}(\lambda k)\underset{\lambda\to 0}{=}\lambda^{2}\mathcal{M}[h_{2};-2]+O\left(\lambda^{3}\right).$ (199)
Keeping the Mellin transform of $h_{2}$ in the integral form, we obtain the expansion for the second term, i.e.,
$\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,f_{2}(k)h_{2}(\lambda k)\underset{\lambda\to 0}{\sim}\frac{\lambda^{2}}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{3}}\arg\frac{1-F(k;0)}{1-F_{2}(k;0)}.$ (200)
Combining (196) with (200) gives the expansion for $I(\lambda;0,1)$
$I(\lambda;0,1)\underset{\lambda\to 0}{=}\mathcal{A}_{0}^{(1)}\lambda+\mathcal{A}_{0}^{(2)}\lambda^{2}+O(\lambda^{3}).$ (201)
Recalling the definition (29) of the moments $\mu_{\ell}(\rho)$, the coefficients can be expressed as
$\mathcal{A}_{0}^{(1)}$ $=-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[\frac{1}{k^{2}}\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right],$ (202)
$\mathcal{A}_{0}^{(2)}$ $=-\frac{1}{12}\frac{\left\langle(M-\alpha t)^{3}\right\rangle}{\left\langle M-\alpha t\right\rangle}-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{3}}\,\mathrm{Im}\left[\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right].$ (203)
Having established the $\lambda\to 0$ behavior, we now turn to the expansion of $I(\lambda;0,1)$ as $\lambda\to\infty$.
Expansion $\lambda\to\infty$
In principle, to obtain the expansion of $I(\lambda;0,1)$ as $\lambda\to\infty$, we could rescale the integration variable $\lambda k\mapsto k$ and apply (164) with the roles of $f$ and $h$ reversed. However, this procedure would generate the logarithmic terms (167). A more convenient strategy is to first extract the logarithmic behavior explicitly by rewriting (189) as
$I(\lambda;0,1)=\frac{1}{2}\log\left[\frac{\mu_{2}(0)}{2}\right]+\log\left[\lambda+k_{+}^{*}(0,1)\right]\\ -\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right]-\frac{1}{\pi}\int_{0}^{\infty}\,\mathrm{Re}\left[\frac{\log\left[1-F(k;0)\right]}{\lambda-\mathrm{i}k}\right],$ (204)
where $k_{+}^{*}(0,1)$ is given by (61). To obtain (204) from (189) we essentially went from the regularized representation (62) of $\phi_{\text{a.c.}}^{-}(\lambda;\rho,s)$ back to the original one (55).
The nontrivial $\lambda$-dependence is now contained in the last term. Separating the real and imaginary parts explicitly we rewrite it as
$\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[\frac{\log[1-F(k;0)]}{\lambda-\mathrm{i}k}\right]=\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{1+(k/\lambda)^{2}}\,\mathrm{Re}\Big[\log\left[1-F(k;0)\right]\Big]\\ -\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\;\frac{(k/\lambda)}{1+(k/\lambda)^{2}}\,\mathrm{Im}\Big[\log\left[1-F(k;0)\right]\Big].$ (205)
The asymptotic behavior of (205) can now be found from (164). However, for our purposes it is enough to extract the leading term, which is already clear from (205). Specifically, the leading order behavior is governed solely by the first term, and we have
$\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[\frac{\log[1-F(k;0)]}{\lambda-\mathrm{i}k}\right]\underset{\lambda\to\infty}{\sim}\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\Big[\log\left[1-F(k;0)\right]\Big].$ (206)
Hence the expansion of $I(\lambda;0,1)$ as $\lambda\to\infty$
$I(\lambda;0,1)\underset{\lambda\to\infty}{=}\log\left[\lambda+k_{+}^{*}(0,1)\right]+\mathcal{A}_{\infty}^{(0)}+\frac{1}{\lambda}\mathcal{A}_{\infty}^{(1)}+O\left(\frac{1}{\lambda^{2}}\right),$ (207)
with
$\mathcal{A}_{\infty}^{(0)}$ $=\frac{1}{2}\log\left[\frac{\mu_{2}(0)}{2}\right]-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right],$ (208)
$\mathcal{A}_{\infty}^{(1)}$ $=-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\Big[\log\left[1-F(k;0)\right]\Big].$ (209)
As a final remark, note that the first term in (207) actually differs depending on whether the system is in survival or absorption regime,
$\log\left[\lambda+k_{+}^{*}(0,1)\right]=\begin{cases}\log\lambda,\quad&\langle M-\alpha t\rangle<0\\[10.0pt] \log\left[\lambda+2\frac{\langle M-\alpha t\rangle}{\langle(M-\alpha t)^{2}\rangle}\right],\quad&\langle M-\alpha t\rangle\geq 0\end{cases}$ (210)
A.2. Expansion of $\partial_{s}I(\lambda;0,s)$
The next step in our asymptotic analysis is to build the expansions for the derivative $\partial_{s}I(\lambda;0,s)|_{s=1}$, which will appear in the representation for the variance of $n$.
From (160) it immediately follows that
$\partial_{s}I(\lambda;0,s)\big|_{s=1}=-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Re}\left[\left(\frac{1}{\mathrm{i}k}+\frac{1}{\lambda-\mathrm{i}k}\right)\left(\frac{F_{2}(k;0)}{1-F_{2}(k,0)}-\frac{F(k;0)}{1-F(k;0)}\right)\right].$ (211)
Separating real and imaginary parts and rescaling the integration variable yields decomposition of (211) into a sum of two Mellin convolution integrals:
$\partial_{s}I(\lambda;0,s)\Big|_{s=1}=-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;f_{1}(k)g_{1}(\lambda k)-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;f_{2}(k)g_{2}(\lambda k),$ (212)
where $f_{1,2}(k)$ are given by (190) and we introduced two new functions given by
$g_{1}(k)$ $=\,\mathrm{Re}\left[\left(\frac{1}{1-F_{2}(k,0)}-\frac{1}{1-F(k;0)}\right)\right],$ (213)
$g_{2}(k)$ $=\,\mathrm{Im}\left[\left(\frac{1}{1-F_{2}(k,0)}-\frac{1}{1-F(k;0)}\right)\right].$
We now apply the general result (164) to each term in (212) separately and obtain $\lambda\to 0$ and $\lambda\to\infty$ expansions.
Expansion $\lambda\to 0$
For the first term in (212), the required behaviors are
$f_{1}(k)\underset{k\to\infty}{=}\frac{1}{k^{2}}+O\left(\frac{1}{k^{4}}\right),\quad g_{1}(k)\underset{k\to 0}{=}O\left(k^{2}\right).$ (214)
Hence in the notation (163) we have
$a_{0}=1,\quad\alpha_{0}=2,\quad\alpha_{1}=4,\quad\beta_{0}=2,$ (215)
and therefore the asymptotic expansion reads
$\int_{0}^{\infty}\mathrm{d}k\;f_{1}(k)g_{1}(\lambda k)\underset{\lambda\to 0}{=}\lambda\;\mathcal{M}[g_{1};-1]+O\left(\lambda^{2}\right),$ (216)
or in the explicit way
$\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;f_{1}(k)g_{1}(\lambda k)\underset{\lambda\to 0}{\sim}\frac{\lambda}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{2}}\,\mathrm{Re}\left[\frac{1}{1-F_{2}(k,0)}-\frac{1}{1-F(k;0)}\right].$ (217)
Similarly for the convoluted functions $f_{2}(k)$ and $g_{2}(k)$ in the second term in (212) the required expansions are
$f_{2}(k)\underset{k\to\infty}{\sim}O\left(\frac{1}{k^{3}}\right),\quad g_{2}(k)\underset{k\to 0}{=}-\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)^{2}}k+O(k^{2}),$ (218)
so that
$\alpha_{0}=3,\quad b_{0}=-\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)^{2}},\quad\beta_{0}=1,\quad\beta_{1}=2,$ (219)
and hence the behavior
$\int_{0}^{\infty}\mathrm{d}k\;f_{2}(k)g_{2}(\lambda k)\underset{\lambda\to 0}{=}-\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)^{2}}\lambda\;\mathcal{M}[f_{2};2]+O(\lambda^{2}).$ (220)
The Mellin transform of $f_{2}(k)$ can be computed explicitly resulting in
$\mathcal{M}[f_{2};z]=\int_{0}^{\infty}\mathrm{d}k\;k^{z-1}\frac{1}{k(k^{2}+1)}=-\frac{\pi}{2}\frac{1}{\cos\frac{\pi z}{2}},\qquad\mathcal{M}[f_{2};2]=\frac{\pi}{2}.$ (221)
Substituting now (221) into (220) we obtain the asymptotic behavior of the second term in (212)
$\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;f_{2}(k)g_{2}(\lambda k)\underset{\lambda\to 0}{\sim}\frac{1}{12}\frac{\mu_{3}(0)}{\mu_{1}(0)^{2}}\lambda$ (222)
Combining (222) with (217) we arrive at the asymptotic expansion of $\partial_{s}I(\lambda;0,s)|_{s=1}$ as $\lambda\to 0$. Specifically, recalling the definition of the moments (29) we obtain
$\partial_{s}I(\lambda;0,s)\Big|_{s=1}\underset{\lambda\to 0}{=}\mathcal{B}_{0}^{(1)}\lambda+O(\lambda^{2})$ (223)
where the coefficient is given by
$\mathcal{B}_{0}^{(1)}=\frac{1}{12}\frac{\langle(M-\alpha t)^{3}\rangle}{\langle M-\alpha t\rangle^{2}}-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{2}}\,\mathrm{Re}\left[\frac{1}{1-F_{2}(k,0)}-\frac{1}{1-F(k;0)}\right].$ (224)
Expansion $\lambda\to\infty$
To find the expansion for $\lambda\to\infty$, we rescale the variable of integration $k\mapsto k/\lambda$ in (212), arriving at
$\partial_{s}I(\lambda;0,s)\Big|_{s=1}=-\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\left(f_{1}(k/\lambda)g_{1}(k)+f_{2}(k/\lambda)g_{2}(k)\right).$ (225)
This is again a sum of two Mellin convolution integrals with roles of the $f$ and $g$ reversed. Now, we shall use the asymptotic behaviors of $f_{1}$ as $k\to 0$ and $g_{1}$ as $k\to\infty$, i.e.,
$g_{1}(k)\underset{k\to\infty}{=}-1+O\left(\frac{1}{k^{2}}\right),\quad f_{1}(k)\underset{k\to 0}{=}1+O(k^{2}).$ (226)
Note that the next term in the expansion of $g_{1}(k)$ may contain oscillating components when the distributions $p(t)$ or $q(M)$ have bounded support, but this affects only the terms in the expansion of (225) of order $O(1/\lambda^{2})$ and higher.
Expansions (226) when combined with the general result (164) imply that
$a_{0}=-1,\quad\alpha_{0}=0,\quad\alpha_{1}=2,\quad b_{0}=1,\quad\beta_{0}=0,\quad\beta_{1}=2,$ (227)
and hence
$\int_{0}^{\infty}\mathrm{d}k\;f_{1}(k/\lambda)g_{1}(k)\underset{\lambda\to\infty}{=}-\lambda\;\mathcal{M}[f_{1};1]+\mathcal{M}[g_{1};1]+O\left(\frac{1}{\lambda}\right).$ (228)
Now we need to compute the Mellin transforms. The Mellin transform of $f_{1}$ can be computed explicitly as in (195), i.e.,
$\mathcal{M}\left[f_{1};z\right]=\int_{0}^{\infty}\mathrm{d}k\;k^{z-1}\frac{1}{k^{2}+1}=\frac{\pi}{2}\frac{1}{\sin\frac{\pi z}{2}},\qquad\mathcal{M}[f_{1};1]=\frac{\pi}{2}.$ (229)
The transform of $g_{1}$ is however more subtle, as the integral (165) diverges, and hence the Mellin transform should be understood in the analytical continuation sense. The procedure is explained in [63]; we state only the result here
$\mathcal{M}[g_{1};1]=\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[1+\left(\frac{1}{1-F_{2}(k,0)}-\frac{1}{1-F(k;0)}\right)\right].$ (230)
Substituting (230) and (229) into (228) after simplification yields
$\int_{0}^{\infty}\mathrm{d}k\;f_{1}(k/\lambda)g_{1}(k)\underset{\lambda\to\infty}{\sim}-\frac{\pi\lambda}{2}+\frac{\pi}{2\left|\mu_{1}(0)\right|}-\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[\frac{F(k;0)}{1-F(k;0)}\right].$ (231)
Similarly, for the second integral in (225) we shall utilize the expansions of $g_{2}(k)$ as $k\to\infty$ and $f_{2}(k)$ as $k\to 0$ that are easily extracted from (190) and (213),
$g_{2}(k)\underset{k\to\infty}{=}O\left(\frac{1}{k^{2}}\right),\quad f_{2}(k)\underset{k\to 0}{=}\frac{1}{k}+O\left(k\right).$ (232)
In the notation of (164) this corresponds to
$\alpha_{0}=2,\quad b_{0}=1,\quad\beta_{0}=-1,\quad\beta_{1}=1,$ (233)
hence the asymptotic expansion
$\int_{0}^{\infty}\mathrm{d}k\;f_{2}(k/\lambda)g_{2}(k)\underset{\lambda\to\infty}{=}\lambda\;\mathcal{M}[g_{2};0]+O\left(\frac{1}{\lambda}\right),$ (234)
or with the explicit form of the Mellin transform
$\int_{0}^{\infty}\mathrm{d}k\;f_{2}(k/\lambda)g_{2}(k)\underset{\lambda\to\infty}{\sim}\lambda\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k}\,\mathrm{Im}\left[\frac{1}{1-F_{2}(k;0)}-\frac{1}{1-F(k;0)}\right].$ (235)
Substituting (231) and (235) into (225) yields the asymptotic expansion of $\partial_{s}I(\lambda;0,s)|_{s=1}$ as $\lambda\to\infty$
$\partial_{s}I(\lambda;0,s)\Big|_{s=1}\underset{\lambda\to\infty}{=}\mathcal{B}_{\infty}^{(0)}+\frac{1}{\lambda}\mathcal{B}_{\infty}^{(1)}+O\left(\frac{1}{\lambda^{2}}\right),$ (236)
with
$\mathcal{B}_{\infty}^{(0)}$ $=\frac{1}{2}-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k}\,\mathrm{Im}\left[\frac{1}{1-F_{2}(k;0)}-\frac{1}{1-F(k;0)}\right],$ (237)
$\mathcal{B}_{\infty}^{(1)}$ $=-\frac{1}{2\left|\left\langle M-\alpha t\right\rangle\right|}+\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[\frac{F(k;0)}{1-F(k;0)}\right].$ (238)
A.3. Expansion of $\partial_{\rho}I(\lambda;\rho,1)$
Finally we proceed to the asymptotic expansions of $\partial_{\rho}I(\lambda;\rho,1)|_{\rho=0}$. Taking the derivative of (160) we arrive at
$\partial_{\rho}I(\lambda;\rho,1)\Big|_{\rho=0}=-\frac{1}{\pi\alpha}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[\frac{1}{(k+\mathrm{i}\lambda)^{2}}\log\frac{1-F(k;\rho)}{1-F_{2}(k;\rho)}\right]\\ -\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[\frac{\lambda}{k(k+\mathrm{i}\lambda)}\left.\left(\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;\rho)}-\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right)\right|_{\rho=0}\right].$ (239)
Separating real and imaginary parts yields a decomposition into a sum of four Mellin convolution integrals, as
$\partial_{\rho}I(\lambda;\rho,1)\Big|_{\rho=0}=-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;f_{1}(k)r_{1}(\lambda k)-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;f_{2}(k)r_{2}(\lambda k)\\ -\frac{1}{\pi\lambda\alpha}\int_{0}^{\infty}\mathrm{d}k\;f_{3}(k)h_{1}(\lambda k)-\frac{1}{\pi\lambda\alpha}\int_{0}^{\infty}\mathrm{d}k\;f_{4}(k)h_{2}(\lambda k)$ (240)
with the functions $f_{1,2}(k)$, $h_{1,2}(k)$ given by (190) and (191), and four additional auxiliary functions defined as
$f_{3}(k)=\frac{k^{2}-1}{\left(k^{2}+1\right)^{2}},\qquad f_{4}(k)=\frac{2k}{\left(k^{2}+1\right)^{2}},$ (241)
$r_{1}(k)$ $=\,\mathrm{Re}\left.\left[\frac{\partial_{\rho}c(\rho)F_{2}(\lambda k;\rho)}{1-F_{2}(\lambda k;\rho)}-\frac{\partial_{\rho}c(\rho)F(\lambda k;\rho)}{1-F(\lambda k;\rho)}\right]\right|_{\rho=0},$ (242)
$r_{2}(k)$ $=\,\mathrm{Im}\left.\left[\frac{\partial_{\rho}c(\rho)F_{2}(\lambda k;\rho)}{1-F_{2}(\lambda k;\rho)}-\frac{\partial_{\rho}c(\rho)F(\lambda k;\rho)}{1-F(\lambda k;\rho)}\right]\right|_{\rho=0}.$
The technique however remains the same, that is we apply (164) to each of four integrals separately.
Expansion $\lambda\to 0$
The functions $r_{1}(k)$ and $r_{2}(k)$ contain the factor $\partial_{\rho}c(\rho)F(k;\rho)$. Let us first take a closer look at it. From (10), (11) it follows that
$\partial_{\rho}c(\rho)F(k;\rho)\Big|_{\rho=0}\underset{k\to 0}{\sim}-\frac{\left\langle M\right\rangle}{\alpha}-\mathrm{i}k\;\frac{\left\langle M(M-\alpha t)\right\rangle}{\alpha}\\ +k^{2}\;\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{2\alpha}+\mathrm{i}k^{3}\;\frac{\left\langle M(M-\alpha t)^{3}\right\rangle}{6\alpha},$ (243)
and similarly from (59) we obtain
$\partial_{\rho}c(\rho)F_{2}(k;\rho)\Big|_{\rho=0}=-\frac{\left\langle M\right\rangle}{\alpha}-\mathrm{i}k\;\frac{\left\langle M(M-\alpha t)\right\rangle}{\alpha}+k^{2}\;\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{2\alpha}.$ (244)
Therefore the relevant asymptotic expansions for the first term are
$f_{1}(k)\underset{k\to\infty}{=}\frac{1}{k^{2}}+O(1),\quad r_{1}(k)\underset{k\to 0}{=}O\left(k^{2}\right)$ (245)
which in notations (164) corresponds to
$a_{0}=1,\quad\alpha_{0}=2,\quad\alpha_{1}=0,\quad\beta_{0}=2,$ (246)
hence the asymptotic behavior
$\int_{0}^{\infty}\mathrm{d}k\,f_{1}(k)r_{1}(\lambda k)=\lambda\,\mathcal{M}\left[r_{1};-1\right]+O(\lambda^{2}).$ (247)
In the explicit form it reads
$\int_{0}^{\infty}f_{1}(k)r_{1}(\lambda k)\underset{\lambda\to 0}{\sim}\lambda\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{2}}\,\mathrm{Re}\left.\left[\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;\rho)}-\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right]\right|_{\rho=0}.$ (248)
For the second term in (212) the asymptotic behaviors are
$f_{2}(k)\underset{k\to\infty}{=}\frac{1}{k^{3}}+O\left(\frac{1}{k}\right),\quad r_{2}(k)\underset{k\to 0}{=}\frac{\left\langle M\right\rangle}{6\alpha}\frac{\mu_{3}(0)}{\mu_{1}(0)^{2}}k+O\left(k^{2}\right)$ (249)
which implies
$a_{0}=1,\quad\alpha_{0}=3,\quad\alpha_{1}=1,\quad b_{0}=\frac{\left\langle M\right\rangle}{6\alpha}\frac{\mu_{3}(0)}{\mu_{1}(0)^{2}},\quad\beta_{0}=1,\quad\beta_{1}=2.$ (250)
This leads to
$\int_{0}^{\infty}\mathrm{d}k\,f_{2}(k)r_{2}(\lambda k)\underset{\lambda\to 0}{=}\frac{\left\langle M\right\rangle}{6\alpha}\frac{\mu_{3}(0)}{\mu_{1}(0)^{2}}\,\lambda\,\mathcal{M}\left[f_{2};2\right]+O(\lambda^{2}).$ (251)
Computing the Mellin transform of $f_{2}$ as in (221) and using the explicit representation for the moments (29) we obtain
$\int_{0}^{\infty}\mathrm{d}k\;f_{2}(k)r_{2}(\lambda k)\underset{\lambda\to 0}{\sim}\lambda\,\frac{\pi}{12}\frac{\left\langle M\right\rangle}{\alpha}\frac{\left\langle(M-\alpha t)^{3}\right\rangle}{\left\langle M-\alpha t\right\rangle^{2}}.$ (252)
The third term in (240) requires the expansions
$f_{3}(k)\underset{k\to\infty}{=}\frac{1}{k^{2}}+O\left(\frac{1}{k^{4}}\right),\quad h_{1}(k)\underset{k\to 0}{=}-\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)}k^{2}+O\left(k^{3}\right)$ (253)
which means
$a_{0}=1,\quad\alpha_{0}=2,\quad\alpha_{1}=4,\quad b_{0}=-\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)},\quad\beta_{0}=2,\quad\beta_{1}=3.$ (254)
Hence the asymptotic expansion
$\int_{0}^{\infty}\mathrm{d}k\,f_{2}(k)h_{1}(\lambda k)\underset{\lambda\to 0}{=}\lambda\,\mathcal{M}[h_{1};-1]-\frac{1}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)}\lambda^{2}\,\mathcal{M}\left[f_{3};3\right]+O\left(\lambda^{3}\right).$ (255)
The Mellin transform of $f_{3}$ can be computed straightforwardly and reads
$\mathcal{M}[f_{3};z]=\int_{0}^{\infty}\mathrm{d}k\,k^{z-1}\frac{k^{2}-1}{(k^{2}+1)^{2}}=\frac{\pi}{2}\frac{z-1}{\sin\frac{\pi}{2}z},\qquad\mathcal{M}[f_{3};3]=-\pi.$ (256)
The asymptotic expansion for the third term now reads
$\int_{0}^{\infty}\mathrm{d}k\,f_{3}(k)h_{1}(\lambda k)\underset{\lambda\to 0}{\sim}\lambda\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{2}}\,\mathrm{Re}\left[\log\frac{1-F(k;0)}{1-F_{2}(k,0)}\right]+\lambda^{2}\,\frac{\pi}{6}\frac{\mu_{3}(0)}{\mu_{1}(0)}.$ (257)
For the last term we have
$f_{4}(k)\underset{k\to\infty}{=}\frac{2}{k^{3}}+O\left(\frac{1}{k^{5}}\right),\quad h_{2}(k)\underset{k\to 0}{=}O(k^{3}),$ (258)
hence
$a_{0}=2,\quad\alpha_{0}=3,\quad\alpha_{1}=5,\quad\beta_{0}=3,$ (259)
and the asymptotic expansion of the integral is
$\int_{0}^{\infty}\mathrm{d}k\,f_{4}(k)h_{2}(\lambda k)\underset{\lambda\to 0}{=}2\lambda^{2}\mathcal{M}[h_{2},-2]+O\left(\lambda^{3}\right),$ (260)
or explicitly
$\int_{0}^{\infty}f_{4}(k)h_{2}(\lambda k)\underset{\lambda\to 0}{\sim}2\lambda^{2}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{3}}\,\mathrm{Im}\left[\log\frac{1-F(k;0)}{1-F_{2}(k,0)}\right].$ (261)
Finally, we substitute (248), (252), (257), and (261) into (240) to obtain the $\lambda\to 0$ expansion of $\partial_{\rho}I(\lambda;\rho,1)|_{\rho=0}$, namely
$\partial_{\rho}I(\lambda;\rho,1)\Big|_{\rho=0}\underset{\lambda\to 0}{=}\mathcal{C}_{0}^{(0)}+\mathcal{C}_{0}^{(1)}\lambda+O(\lambda^{2}),$ (262)
where the coefficients are given by
$\mathcal{C}_{0}^{(0)}$ $=-\frac{1}{\pi\alpha}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k^{2}}\,\mathrm{Re}\left[\log\frac{1-F(k;0)}{1-F_{2}(k,0)}\right]=\frac{1}{\alpha}\mathcal{A}_{0}^{(1)}$ (263)
$\mathcal{C}_{0}^{(1)}$ $=-\frac{1}{6\alpha}\frac{\langle(M-\alpha t)^{3}\rangle}{\langle M-\alpha t\rangle}\left(1+\frac{\left\langle M\right\rangle}{2\langle M-\alpha t\rangle}\right)$
$\qquad-\frac{2}{\alpha\pi}\int_{0}^{\infty}\mathrm{d}k\,\frac{1}{k^{3}}\,\mathrm{Im}\left[\log\frac{1-F(k;0)}{1-F_{2}(k,0)}\right]$
$\qquad-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\frac{1}{k^{2}}\,\mathrm{Re}\left.\left[{}\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;\rho)}-\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right]\right|_{\rho=0}.$ (264)
Expansion $\lambda\to\infty$
To shorten the calculation a little bit, let us look at the second term in (239). Using an integral identity
$\int_{0}^{\infty}\mathrm{d}k\;\,\mathrm{Re}\left[\frac{\log\left[1-F_{2}(k;\rho)\right]}{(k+\mathrm{i}\lambda)^{2}}\right]=\frac{\pi}{\lambda}$ (265)
which can be verified by direct calculation we can simplify it as
$\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Re}\left[\frac{1}{(k+\mathrm{i}\lambda)^{2}}\log\frac{1-F(k;\rho)}{1-F_{2}(k;\rho)}\right]=-\frac{\pi}{\lambda}\\ +\frac{1}{\lambda^{2}}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Re}\left[\frac{\log[1-F(k;\rho)]}{\left((k/\lambda)^{2}+\mathrm{i}\right)^{2}}\right].$ (266)
The second term impacts the expansion as $\lambda\to\infty$ only at the order higher than $\lambda^{-2}$. Therefore after rescaling (240) transforms into
$\partial_{\rho}I(\lambda;\rho,1)\Big|_{\rho=0}\underset{\lambda\to\infty}{=}\frac{1}{\alpha\lambda}-\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}kf_{1}\left(k/\lambda\right)r_{1}(k)\\ -\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}kf_{2}\left(k/\lambda\right)r_{2}(k)+O\left(\frac{1}{\lambda^{2}}\right),$ (267)
where the functions $f_{1,2}(k)$ and $r_{1,2}(k)$ are given by (190) and (242) respectively. Thus, we need to apply (164) only to two terms out of four in (240).
Now we need the expansions of $r_{1,2}(k)$ for $k\to\infty$. A simple calculation shows that due to (33) we have
$\left.\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right|_{\rho=0}\underset{k\to\infty}{=}O\left(\frac{1}{k^{2}}\right),$ (268)
and using the explicit form (59) of $F_{2}(k;\rho)$ we obtain
$\left.\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right|_{\rho=0}\underset{k\to\infty}{=}-\frac{1}{\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}\\ +\frac{2\mathrm{i}}{k}\left(\frac{\left\langle M(M-\alpha t)\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}-\frac{\left\langle M-\alpha t\right\rangle\left\langle M(M-\alpha t)^{2}\right\rangle}{\big[\left\langle(M-\alpha t)^{2}\right\rangle\big]^{2}}\right)+O\left(\frac{1}{k^{2}}\right).$ (269)
Therefore, only the term involving $F_{2}(k;\rho)$ contributes to the asymptotic behavior of $r_{1,2}(k)$ for $k\to\infty$. Note, however, that this does not mean that $\lambda\to\infty$ expansion does not depend on $F(k;\rho)$.
Now we apply (164) to the first integral term in (267). The relevant expansions are
$r_{1}(k)\underset{k\to\infty}{=}\frac{1}{\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}+O\left(\frac{1}{k^{2}}\right),\qquad f_{1}(k)\underset{k\to 0}{=}1+O\left(k^{2}\right),$ (270)
hence the coefficients in (164)
$a_{0}=\frac{1}{\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle},\quad\alpha_{0}=0,\quad\alpha_{1}=1,b_{0}=1,\quad\beta_{0}=0,\quad\beta_{1}=2$ (271)
which implies
$\int_{0}^{\infty}\mathrm{d}kf_{1}\left(k/\lambda\right)r_{1}(k)\underset{\lambda\to\infty}{=}\frac{1}{\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}\,\lambda\,\mathcal{M}[f_{1};1]+\mathcal{M}\left[r_{1};1\right]+O\left(\frac{1}{\lambda}\right).$ (272)
The Mellin transform of $f_{1}$ we computed in (195), and it reads
$\mathcal{M}\left[f_{1};z\right]=\int_{0}^{\infty}\mathrm{d}k\;k^{z-1}\frac{1}{k^{2}+1}=\frac{\pi}{2}\frac{1}{\sin\frac{\pi z}{2}},\qquad\mathcal{M}[f_{1};1]=\frac{\pi}{2}.$ (273)
As for the Mellin transform of $r_{1}$, we face the same problem as for $g_{1}$ in (230). Specifically, the integral (165) diverges and we need to perform the analytic continuation as prescribed in [63]. The resulting expression is
$\mathcal{M}[r_{1};1]=\int_{0}^{\infty}\mathrm{d}k\,\,\mathrm{Re}\left.\left[\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;\rho)}-\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}-\frac{1}{\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}\right]\right|_{\rho=0}.$ (274)
This integral can be simplified if we compute the contribution of the term with $F_{2}(k,\rho)$, specifically
$\int_{0}^{\infty}\mathrm{d}k\,\,\mathrm{Re}\left.\left[\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;\rho)}-\frac{1}{\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}\right]\right|_{\rho=0}\\ =\frac{\pi}{2\alpha}\frac{\left\langle M\right\rangle}{\left\langle M-\alpha t\right\rangle}+\frac{\pi}{\alpha}\frac{\left\langle M-\alpha t\right\rangle\left\langle M(M-\alpha t)^{2}\right\rangle}{[\left\langle(M-\alpha t)^{2}\right\rangle]^{2}}-\frac{\pi}{\alpha}\frac{\left\langle M(M-\alpha t)\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle},$ (275)
and hence the expansion
$\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\,f_{1}(k/\lambda)r_{1}(k)\underset{\lambda\to\infty}{\sim}-\frac{1}{2\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}\\ -\frac{1}{\alpha\lambda}\left\{\frac{1}{2}\frac{\left\langle M\right\rangle}{\left\langle M-\alpha t\right\rangle}+\frac{\left\langle M-\alpha t\right\rangle\left\langle M(M-\alpha t)^{2}\right\rangle}{[\left\langle(M-\alpha t)^{2}\right\rangle]^{2}}-\frac{\left\langle M(M-\alpha t)\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}\right\}\\ +\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\,\,\mathrm{Re}\left.\left[\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right]\right|_{\rho=0}.$ (276)
Finally, for the second integral term in (267) the relevant expansions are
$r_{2}(k)\underset{k\to\infty}{=}\frac{a_{0}}{k}+O\left(\frac{1}{k^{3}}\right),\quad f_{2}(k)\underset{k\to 0}{=}\frac{1}{k}+O\left(k\right),$ (277)
where
$a_{0}=2\frac{\left\langle M-\alpha t\right\rangle\left\langle M(M-\alpha t)^{2}\right\rangle}{\big[\left\langle(M-\alpha t)^{2}\right\rangle\big]^{2}}-2\frac{\left\langle M(M-\alpha t)\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}.$ (278)
Then according to (164) we have
$\alpha_{0}=1,\quad\alpha_{1}=3,\quad b_{0}=1,\quad\beta_{0}=-1,\quad\beta_{1}=1,$ (279)
hence
$\int_{0}^{\infty}\mathrm{d}kf_{2}\left(k/\lambda\right)r_{2}(k)\underset{\lambda\to\infty}{=}\lambda\mathcal{M}[r_{2};0]+a_{0}\mathcal{M}\left[f_{2};0\right].$ (280)
The Mellin transform of $f_{2}$ is given by (221) and it reads
$\mathcal{M}[f_{2};z]=\int_{0}^{\infty}\mathrm{d}k\;k^{z-1}\frac{1}{k(k^{2}+1)}=-\frac{\pi}{2}\frac{1}{\cos\frac{\pi z}{2}},\qquad\mathcal{M}[f_{2};0]=-\frac{\pi}{2},$ (281)
hence the expansion
$\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\,f_{2}(k/\lambda)r_{2}(k)\underset{\lambda\to\infty}{\sim}\\ -\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k}\,\mathrm{Im}\left.\left[\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;\rho)}-\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right]\right|_{\rho=0}\\ +\frac{1}{\lambda}\left\{\frac{\left\langle M-\alpha t\right\rangle\left\langle M(M-\alpha t)^{2}\right\rangle}{[\left\langle(M-\alpha t)^{2}\right\rangle]^{2}}-\frac{\left\langle M(M-\alpha t)\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}\right\}.$ (282)
Combining now (282) and (276) with (267) we arrive at the expansion
$\partial_{\rho}I(\lambda;\rho,1)\Big|_{\rho=0}\underset{\lambda\to\infty}{=}\mathcal{C}_{\infty}^{(0)}+\frac{1}{\lambda}\mathcal{C}_{\infty}^{(1)}+O\left(\frac{1}{\lambda^{2}}\right)$ (283)
with
$\mathcal{C}_{\infty}^{(0)}$ $=-\frac{1}{2\alpha}\frac{\left\langle M(M-\alpha t)^{2}\right\rangle}{\left\langle(M-\alpha t)^{2}\right\rangle}$
$\quad-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\frac{1}{k}\,\mathrm{Im}\left.\left[\frac{\partial_{\rho}c(\rho)F_{2}(k;\rho)}{1-F_{2}(k;\rho)}-\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right]\right|_{\rho=0}$ (284)
$\mathcal{C}_{\infty}^{(1)}$ $=\frac{1}{\alpha}-\frac{1}{2\alpha}\frac{\left\langle M\right\rangle}{\left\langle M-\alpha t\right\rangle}+\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\,\mathrm{Re}\left.\left[\frac{\partial_{\rho}c(\rho)F(k;\rho)}{1-F(k;\rho)}\right]\right|_{\rho=0}.$ (285)
Note that in (284), we can actually extract the term with $F_{2}(k;\rho)$ by explicit integration. However, this provides no simplification, so we keep (284) in the present form.
A.4. Asymptotic expansion for the moments
Having computed all necessary expansions of $I(\lambda;0,s)$ and its first two derivatives, we now proceed to constructing asymptotic expansions of the mean and variances of $\tau$ and $n$. We provide the detailed derivation for the mean of $n$ as an illustration. The variance of $n$ and the moments of $\tau$ follow by the same procedure, and we present only the main steps.
Mean values
Taking the limit $\rho\to 0$ in (64) and expanding it in series we find that
$\frac{\partial}{\partial s}\hat{Q}(0,s\,|\,\lambda)\Big|_{s=1}=-\frac{1}{\lambda^{2}}\frac{e^{I(\lambda;0,1)}}{\left\langle M-\alpha t\right\rangle}.$ (286)
For the expansion as $\lambda\to 0$, we substitute (201) into (286) to obtain
$\frac{\partial}{\partial s}\hat{Q}(0,s\,|\,\lambda)\Big|_{s=1}\underset{\lambda\to 0}{=}-\frac{1}{\lambda^{2}}\frac{1}{\left\langle M-\alpha t\right\rangle}-\frac{1}{\lambda}\frac{\mathcal{A}_{0}^{(1)}}{\left\langle M-\alpha t\right\rangle}+O(1).$ (287)
This suggests a linear behavior of the mean value, namely
$\mathbb{E}[n\,|\,X_{0}]\underset{X_{0}\to\infty}{\sim}A_{1}X_{0}+A_{0}.$ (288)
Comparing the Laplace transform of the ansatz (288) with (287), we fix the coefficients as
$A_{1}=-\frac{1}{\langle M-\alpha t\rangle},\qquad A_{0}=-\frac{\mathcal{A}_{0}^{(1)}}{\langle M-\alpha t\rangle}.$ (289)
Recalling the definition (202) of $\mathcal{A}_{0}^{(1)}$, we immediately obtain the representation (181).
For the expansion as $\lambda\to\infty$, we substitute (207) into (286). In the absorption regime, $\log[\lambda+k_{+}^{*}(0,1)]=\log\lambda$. Expanding the exponential then yields
$\left.\frac{\partial}{\partial s}\hat{Q}(0,s\,|\,\lambda)\right|_{s=1}\underset{\lambda\to\infty}{=}-\frac{e^{\mathcal{A}_{\infty}^{(0)}}}{\left\langle M-\alpha t\right\rangle}\left(\frac{1}{\lambda}+\frac{\mathcal{A}_{\infty}^{(1)}}{\lambda^{2}}\right)+O\left(\frac{1}{\lambda^{3}}\right).$ (290)
This suggests the behavior
$\mathbb{E}[n\,|\,X_{0}]\underset{X_{0}\to 0}{\sim}a_{0}+a_{1}X_{0},$ (291)
with coefficients
$a_{0}=-\frac{e^{\mathcal{A}_{\infty}^{(0)}}}{\left\langle M-\alpha t\right\rangle},\qquad a_{1}=-\frac{e^{\mathcal{A}_{\infty}^{(0)}}}{\left\langle M-\alpha t\right\rangle}\mathcal{A}_{\infty}^{(1)}.$ (292)
Using the explicit forms of $\mathcal{A}_{\infty}^{(0)}$ and $\mathcal{A}_{\infty}^{(1)}$ given in (208) and (209), we recover the coefficients (182). This completes the analysis for the mean of $n$.
The asymptotic behavior of the mean value of $\tau$ for both $X_{0}\to 0$ and $X_{0}\to\infty$ is readily restored from (288) and (291) since due to (159) the mean values of $\tau$ and $n$ are related via
$\mathbb{E}\left[\tau\,|\,X_{0}\right]=\frac{X_{0}}{\alpha}+\frac{\left\langle M\right\rangle}{\alpha}\mathbb{E}\left[n\,|\,X_{0}\right].$ (293)
Variances
To compute the variance of $n$ we first construct the asymptotic expansion for the second moment, and then combine the result with (288) and (291) we have obtained for the mean value.
The Laplace transform of the second moment in terms of $\hat{Q}(\lambda;\rho,s)$ is given by
$\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\mathbb{E}\left[n^{2}\,|\,X_{0}\right]=\left.\left[\frac{\partial^{2}}{\partial s^{2}}+\frac{\partial}{\partial s}\right]\hat{Q}(0,s\,|\,\lambda)\right|_{s=1}.$ (294)
Therefore we need to compute the asymptotic expansions of the second derivative. A straightforward computation yields
$\frac{\partial^{2}}{\partial s^{2}}\hat{Q}(0,s\,|\,\lambda)\Big|_{s=1}=\frac{2}{\lambda^{3}}\frac{e^{I(\lambda;0,1)}}{\left\langle M-\alpha t\right\rangle^{2}}\\ +\frac{1}{\lambda^{2}}\frac{e^{I(\lambda;0,1)}}{\left\langle M-\alpha t\right\rangle}\left[\frac{\left\langle(M-\alpha t)^{2}\right\rangle}{\left\langle M-\alpha t\right\rangle^{2}}+2\left(\frac{\partial}{\partial s}I(\lambda;0,s)\Big|_{s=1}-1\right)\right].$ (295)
The behavior as $X_{0}\to\infty$ is governed by the $\lambda\to 0$ expansion of (295), which can be constructed straightforwardly using the expansions (201) for $I(\lambda;0,1)$ and (223) for $\partial_{s}I(\lambda;0,s)|_{s=1}$. Substituting the resulting expansions into (294), we find quadratic behavior for the second moment in the form:
$\mathbb{E}\left[n^{2}\,|\,X_{0}\right]\underset{X_{0}\to\infty}{\sim}\sigma_{2}X_{0}^{2}+\sigma_{1}X_{0}+\sigma_{0},$ (296)
where the coefficients $\sigma_{i}$ are represented in terms of $\mathcal{A}_{0}^{(1)}$, $\mathcal{A}_{0}^{(2)}$ and $\mathcal{B}_{0}^{(1)}$ given by (202), (203) and (224). Then using the definition of the variance,
$\mathrm{Var}\left[n\,|\,X_{0}\right]=\mathbb{E}\left[n^{2}\,|\,X_{0}\right]-\big(\mathbb{E}\left[n\,|\,X_{0}\right]\big)^{2},$ (297)
we obtain the asymptotic behavior
$\mathrm{Var}\left[n\,|\,X_{0}\right]\underset{X_{0}\to\infty}{\sim}B_{1}X_{0}+B_{0},$ (298)
with coefficients (183).
Repeating the same procedure for $\lambda\to\infty$, substituting the expansions (207) and (236) into (295), after elementary algebraic manipulations we reconstruct the behavior of the variance for $X_{0}\to 0$, namely
$\mathrm{Var}\left[n\,|\,X_{0}\right]\underset{X_{0}\to 0}{\sim}b_{0}+b_{1}X_{0},$ (299)
with the coefficients (184).
The variance of $\tau$ is computed in the same way. We start with the Laplace transform of the second moment
$\int_{0}^{\infty}\mathrm{d}X_{0}\,e^{-\lambda X_{0}}\mathbb{E}\left[\tau^{2}\,|\,X_{0}\right]=\left.\frac{\partial^{2}}{\partial\rho^{2}}\hat{Q}(\rho,1\,|\,\lambda)\right|_{\rho=0},$ (300)
which is given by
$\left.\frac{\partial^{2}}{\partial\rho^{2}}\hat{Q}(\rho,1\,|\,\lambda)\right|_{\rho=0}=\frac{2}{\alpha^{2}\lambda^{3}}\left(1-\frac{\langle M\rangle}{\langle M-\alpha t\rangle}\left[1-\frac{\langle M\rangle}{\langle M-\alpha t\rangle}\right]e^{I(\lambda;0,1)}\right)\\ +\frac{2\,\langle M\rangle}{\alpha^{2}\lambda^{2}}\left(\frac{2\langle M(M-\alpha t)\rangle}{\langle M-\alpha t\rangle^{2}}-\langle M\rangle\frac{\langle(M-\alpha t)^{2}\rangle}{\langle M-\alpha t\rangle^{3}}\right)e^{I(\lambda;0,1)}\\ +\frac{2}{\alpha^{2}\lambda^{2}}\frac{\langle M\rangle}{\langle M-\alpha t\rangle}\left(\left.\frac{\partial}{\partial\rho}I(\lambda;\rho,1)\right|_{\rho=0}-\frac{\langle M^{2}\rangle}{\langle M\rangle}\right)e^{I(\lambda;0,1)}.$ (301)
The $X_{0}\to\infty$ behavior is obtained by substituting (201) and (262) into (301), which yields a quadratic form for the second moment. Combining this with the mean value from (288) and the relation (159) between $\tau$ and $n$, we find
$\mathrm{Var}\left[\tau\,|\,X_{0}\right]\underset{X_{0}\to\infty}{\sim}C_{1}X_{0}+C_{0},$ (302)
with coefficients given by (185).
Similarly, for the $X_{0}\to 0$ behavior, we substitute (207) and (283) into (301). Following the same procedure, we obtain
$\mathrm{Var}\left[\tau\,|\,X_{0}\right]\underset{X_{0}\to 0}{\sim}c_{0}+c_{1}X_{0},$ (303)
with coefficients (186).